### v2

# Universal Acceleration of Quantum Relaxation by Reset-Induced Mpemba Effect

Abstract
Slow relaxation processes spanning widely separated timescales pose fundamental challenges for probing steady-state properties and engineering functional quantum systems, such as quantum heat engines and quantum computing devices. We introduce a universal quantum reset protocol that enables exponential acceleration of relaxation in general Markovian open quantum systems. The protocol induces a quantum Mpemba effect, wherein systems initialized farther from equilibrium can relax faster than those closer to it, and crucially, this acceleration persists even when the slowest decaying Lindbladian modes form complex-conjugate pairs. Unlike previous approaches, which typically target a single mode, our protocol eliminates multiple relaxation modes simultaneously. This framework provides a versatile and experimentally feasible tool for controlling quantum relaxation timescales, with broad implications for quantum thermodynamics, computation, and state preparation.
Introduction.—Relaxation processes in open quantum systems, wherein a driven or interacting system settles into an equilibrium or a nonequilibrium steady state, are of fundamental importance in nonequilibrium physics and for practical quantum technologies. The timescales of such relaxations often determine the feasibility and performance of quantum devices. For example, in quantum thermodynamics, the power output of heat engines or transport junctions is limited by how rapidly the working substance reaches its steady state in each cycle [1, 2, 3]. In addition, fast relaxation facilitates ground state laser cooling [4, 5, 6] and the reliable preparation of quantum states [7]. Faster relaxation can also lead to more efficient quantum algorithms [8], particularly in dissipative computational architectures where the output is encoded in the system’s stationary state [8, 9]. Slow relaxation inevitably allows other unwanted dissipative dynamics to consume substantial time resources, compromising the efficiency of the final stabilized state [10, 11]. In contrast, slow relaxation is sometimes desirable—such as when the computational task targets metastable states [12, 13, 14, 15], whose prolonged lifetimes are essential for robust information storage or approximate optimization.
Researchers have developed various techniques to accelerate relaxation toward both thermal equilibrium [16] and nonequilibrium steady states [17, 18, 11]. Most of these methods, however, are restricted to specific systems and rely on intricate external controls. For instance, some approaches require auxiliary Hamiltonians or full knowledge of the system Hamiltonian, limiting their applicability to simple setups. A more general and elegant strategy is to accelerate relaxation by tailoring the initial state distribution, without the need for continuous external driving. A particularly intriguing example of this approach is inspired by the so-called Mpemba effect [19], a counterintuitive relaxation anomaly in which a hotter system can cool faster than a colder one. This phenomenon implies the existence of an optimal initial state that minimizes the relaxation timescale. Theoretical frameworks for identifying such optimal states have been proposed in both classical [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31] and quantum systems [10, 32, 33, 34, 35, 36, 37, 38, 39], see also [40, 41] for recent reviews. However, existing quantum implementations typically impose unrealistic constraints, such as assuming pure initial states or requiring the second-largest eigenvalue of the Lindbladian to be real, which limit their applicability in general settings. Thus, a general and practical method for inducing faster relaxation toward arbitrary steady states in open quantum systems remains elusive. Likewise, a systematic recipe for prolonging the lifetime of metastable states, i.e., for slowing down the relaxation process is also lacking.
In this study, we address these challenges by proposing an experimentally feasible protocol that can both accelerate and decelerate general quantum relaxation processes. The protocol is based on the quantum reset operation [42, 43, 44, 45, 46], which is a key component of many quantum computation tasks [47]. For instance, qubit initialization to the computational ground state $|00...0\rangle$ is typically realized through a reset procedure. Recent experimental advancements have demonstrated the practical implementation of quantum reset, e.g. Aamir et al. have realized an autonomous quantum absorption refrigerator using superconducting circuits, capable of resetting a transmon qubit [48]. Other reset schemes have also been proposed [49, 50, 51, 52]. In our protocol, we introduce a finite-duration reset phase into the open-system dynamics by temporarily coupling the system to a reset channel. We show that this operation can exponentially accelerate relaxation, even when starting from mixed initial states and in the presence of complex decay modes. Remarkably, our method allows for the simultaneous elimination of multiple relaxation modes—a capability absent in previous approaches. Moreover, by applying the protocol selectively, the lifetime of metastable states can be extended, offering a systematic route to suppress relaxation.
Setup.—We consider a general open quantum system defined on a Hilbert space $\mathcal{H}$ of dimension $d$, whose dynamics in the interaction picture are governed by the Lindblad-Gorini-Kossakowski-Sudarshan (LGKS) master equation $d\rho/dt=\mathcal{L}(\rho),$ where the Lindbladian superoperator $\mathcal{L}$ is given by
$\mathcal{L}(\rho)\equiv-i[H,\rho]+\sum_{i}\left[J_{i}\rho J_{i}^{\dagger}-\frac{1}{2}\{J_{i}^{\dagger}J_{i},\rho\}\right].$ (1)
[…]
$\text{Tr}(L_{k}^{\dagger}R_{h})=\delta_{kh}.$ (2)
The dual superoperator $\mathcal{L}^{\dagger}$ governs the Heisenberg-picture dynamics of observables:
$\mathcal{L}^{\dagger}(O)=i[H,O]+\sum_{i}\left[J_{i}OJ_{i}^{\dagger}-\frac{1}{2}\{J_{i}^{\dagger}J_{i},O\}\right].$
Given an initial state $\rho_{0}$, the system state at time $t$ evolves as
$\rho(t)=e^{t\mathcal{L}}[\rho_{0}]=\rho_{ss}+\sum_{k=2}^{d^{2}}e^{\lambda_{k}t}c_{k}R_{k},$ (3)
where $c_{k}\equiv\text{Tr}(L_{k}^{\dagger}\rho_{0})$. The unique stationary state of the open quantum system $\rho_{ss}$ is given by $\rho_{ss}=\lim_{t\rightarrow\infty}\rho(t)=R_{1}$, assuming $\lambda_{1}=0$ is non-degenerate. In the long-time limit, the relaxation is dominated by the slowest decaying mode, and the deviation from stationarity obeys
$||\rho(t)-\rho_{ss}||\sim\exp\left(|\text{Re}\lambda_{2}|t\right),$
so the characteristic relaxation timescale is given by $\tau_{{\rm{\rm rel}}}\sim\frac{1}{|\text{Re}(\lambda_{2})|}$. For clarity of exposition, we first consider the simple case where the second eigenvalue $\lambda_{2}$ is real and non-degenerate, as in [10]. Later, we will generalize the discussion to include cases where the subleading eigenvalues form complex conjugate pairs.
Reset protocol to accelerate relaxation processes.—Recognizing that the slowest decaying mode governs the relaxation timescale, we propose a quantum reset protocol that suppresses its excitation from a general initial state via a finite-duration reset phase, thereby enabling exponential acceleration of relaxation. Specifically, we let the system evolve under modified dynamics for a finite time $t_{s}$, during which it is stochastically reset to a chosen state $|\delta\rangle$ at random times, with events following a Poisson process of rate $r$. The reset state may also be taken as a mixed state $\rho_{\delta}$. After time $t_{s}$, the reset channel is turned off, and the system resumes evolving under the original Lindbladian $\mathcal{L}$. The modified dynamics during the reset phase is governed by a new Lindbladian $\mathcal{L}_{r}$, defined as
$\mathcal{L}_{r}(\rho)=\mathcal{L}(\rho)+r\text{Tr}(\rho)|\delta\rangle\langle\delta|-r\rho.$ (4)
This can be physically implemented by adding reset jump operators [42] $J_{i}^{r}=\sqrt{r}|\delta\rangle\langle\phi_{i}|$ to the Lindbladian (1), where $|\phi_{i}\rangle$ form an auxiliary complete orthonormal basis. When the reset state is a mixed state $\rho_{\delta}$, the new Lindbladian reads $\mathcal{L}_{r}(\rho)=\mathcal{L}(\rho)+r\text{Tr}(\rho)\rho_{\delta}-r\rho$. Importantly, this protocol modifies the relaxation spectrum without altering the eigenmodes. This is because the term $r\text{Tr}(\rho)|\delta\rangle\langle\delta|$ in the modified Lindbladian contributes only to the stationary mode and vanishes on all traceless eigenmodes $R_{i}$ with $i\geq 2$, which satisfy $\mathrm{Tr}(R_{i})=0$ due to $\mathrm{Tr}(L_{1}^{\dagger}R_{i})=0$ and $L_{1}^{\dagger}=\mathbb{I}$ [cf. (2)]. As a result, the modified Lindbladian acts as
$\mathcal{L}_{r}(R_{i})=(\lambda_{i}-r)R_{i},\ i\in\{2,...,d^{2}\},$ (5)
leading to a uniform spectral shift by $-r$ while preserving the mode structure. Let $\rho^{r}(t):=e^{t\mathcal{L}_{r}}\rho_{0}$ be the state at time $t$ under reset dynamics. Its spectral decomposition reads:
$\rho^{r}(t)=$ $e^{t\mathcal{L}_{r}}\rho_{ss}+\sum_{k=2}^{d^{2}}\text{Tr}(L_{k}^{\dagger}\rho_{0})e^{(\lambda_{k}-r)t}R_{k}$
::=:= $\rho_{ss}+\sum_{k}c_{k}^{r}(t)e^{\lambda_{k}t}R_{k},$ (6)
where $c_{k}^{r}(t)\propto\text{Tr[$L_{k}^{\dagger}$$\rho^{r}(t)$]}$ are time-dependent mode amplitudes. Explicitly, these modified coefficients are given by
$c_{k}^{r}(t)\equiv\left[c_{k}-\frac{r\cdot d_{k}}{r-\lambda_{k}}\right]e^{-rt}+\frac{r\cdot d_{k}}{r-\lambda_{k}}e^{-\lambda_{k}t},$ (7)
where $d_{k}=\langle\delta|L_{k}^{\dagger}|\delta\rangle$. By properly choosing $t_{s}$, we can set $c_{2}^{r}(t_{s})=0$, thereby eliminating the overlap with the slowest decaying mode. Solving $c_{2}^{r}(t_{s})=0$ yields
$t_{s}(r)=\frac{1}{r-\lambda_{2}}\ln\left[1-\frac{c_{2}(r-\lambda_{2})}{r\cdot d_{2}}\right].$ (8)
This condition is always achievable by an appropriate choice of reset state $|\delta\rangle$ ensuring $t_{s}(r)\geq 0$. After the reset phase ($t>t_{s}$), the system returns to original Lindbladian evolution, now without the contribution from the slowest mode. The full evolution reads:
$\rho^{r}(t)=$ $\rho_{ss}+\sum_{k}c_{k}^{r}(t)e^{\lambda_{k}t}\Theta(t_{s}-t)R_{k}$
$+\sum_{k}c_{k}^{r}(t_{s})e^{\lambda_{k}t}\Theta(t-t_{s})R_{k},$ (9)
where $\Theta(x)$ is the Heavside step function. This protocol reduces the relaxation time to
$\tau_{{\rm rel}}\sim t_{s}+\frac{1}{|{\rm Re}(\lambda_{3})|},$ (10)
which is generically shorter than the unperturbed timescale $\tau_{{\rm{\rm rel}}}\sim\frac{1}{|\text{Re}(\lambda_{2})|}$.
Our protocol also enables controlled deceleration of relaxation. By choosing $|\delta\rangle$ and $t_{s}$ to amplify the magnitude of $c_{k}^{r}(t)$, one can extend the lifetime of selected metastable modes. This feature is especially useful in quantum metrology and computation, where long-lived metastable states play a functional role. Our reset protocol thus provides a general framework for controlling relaxation processes in open quantum systems.
Extension to complex-conjugate relaxation modes.—We now consider the more general case where the dominant relaxation modes form a complex-conjugate pair. In this case, the long-time dynamics of the density matrix takes the form
$\rho(t)=\rho_{ss}+e^{\lambda_{2}t}c_{2}R_{2}+e^{\lambda_{2}^{\dagger}t}c_{2}^{*}R_{2}^{\dagger}+...,$ (11)
where $c_{2}=\text{Tr}(L_{2}^{\dagger}\rho_{0})$ and $c_{2}^{*}=\text{Tr}(L_{2}\rho_{0})$ is its complex conjugate. Most known strategies are unable to simultaneously suppress both components of such complex relaxation modes.
Our reset protocol, however, naturally addresses this limitation. The modified coefficients under quantum reset are given by
$c_{2}^{r}(t)=\left[c_{2}-\frac{r\cdot d_{2}}{r-\lambda_{2}}\right]e^{-rt}+\frac{r\cdot d_{2}}{r-\lambda_{2}}e^{-\lambda_{2}t}$ ,
$c_{2}^{r,*}(t)=\left[c_{2}^{*}-\frac{r\cdot d_{2}^{*}}{r-\lambda_{2}^{*}}\right]e^{-rt}+\frac{r\cdot d_{2}^{*}}{r-\lambda_{2}^{*}}e^{-\lambda_{2}^{*}t}$
Importantly, we observe that these coefficients remain complex conjugates of each other for all $t$, i.e., $[c_{2}^{r}(t)]^{*}=c_{2}^{r,*}(t)$. Therefore, eliminating the entire complex mode reduces to a single complex condition:
$c_{2}^{r}(t_{s})=0\Leftrightarrow\text{Re}[c_{2}^{r}(t_{s})]=\text{Im}[c_{2}^{r}(t_{s})]=0,$ (12)
because $c_{2}^{r}(t_{s})=0\Rightarrow c_{2}^{r,*}(t_{s})=[c_{2}^{r}(t_{s})]^{*}=0$. This condition can be fulfilled by tuning the reset rate $r$ and the reset duration $t_{s}$, thereby simultaneously eliminating both components of the dominant complex relaxation mode. In practice, it suffices to reduce $|c_{2}^{r}(t_{s})|$ to an exponentially small value ($\text{Re}[c_{2}^{r}(t_{s})]\ll 1,\ \text{Im}[c_{2}^{r}(t_{s})]\ll 1$), which still leads to exponential acceleration of relaxation. We demonstrate this behavior concretely in the following example.
Example.—To illustrate our protocol, we consider a minimal example: a two-level quantum system, which may represent a qubit or a spin-$1/2$ particle. We set the ground state energy $E_{0}=0$ and excited state energy $E_{1}=E$, such that the system Hamiltonian reads
$H=E|1\rangle\langle 1|+\Omega(|0\rangle\langle 1|+|1\rangle\langle 0|)\text{,}$
where $\Omega$ is the intrinsic coherent coupling between the two levels. The system is coupled to a thermal environment through the Lindblad jump operators $J_{0}=|1\rangle\langle 0|=\sigma_{+}$ and $J_{1}=|0\rangle\langle 1|=\sigma_{-}$. The resulting open-system dynamics are governed by the master equation:
$\frac{d\rho(t)}{dt}=$ $-i[H,\rho]+\Gamma_{0}\left(J_{0}\rho J_{1}-\frac{1}{2}\{J_{0}J_{1},\rho\}\right)$
$+\Gamma_{1}\left(J_{1}\rho J_{0}-\frac{1}{2}\{J_{1}J_{0},\rho\}\right),$ (13)
where $\Gamma_{0,1}$ are the transition rates from state $|0\rangle$ to state $|1\rangle$ or converse. Assuming the rates satisfy detailed balance with inverse temperature $\Gamma_{0}e^{-\beta_{\text{env}}E_{0}}=\Gamma_{1}e^{-\beta_{\text{env}}E_{1}}$, i.e., $\Gamma_{0}=\Gamma_{1}e^{-\beta_{\text{env}}E},$ the system evolves to a unique equilibrium state $\rho_{eq}=\frac{e^{-\beta_{\text{env}}H}}{\text{Tr}(e^{-\beta_{\text{env}}H})}$ irrespective of the initial condition. We initialize the system in a general mixed state:
$\hat{\rho}_{0}$ $=\left(\begin{array}[]{cc}\frac{1}{1+e^{-\beta_{0}E}}&ke^{i\phi}\\ ke^{-i\phi}&\frac{e^{-\beta_{0}E}}{1+e^{-\beta_{0}E}}\end{array}\right),$
$0\leq k$ $\leq\frac{\sqrt{e^{\beta_{0}E}}}{e^{\beta_{0}E}+1},\quad 0\leq\phi\leq 2\pi,$
where the off-diagonal terms $ke^{\pm i\phi}$ quantify initial quantum coherence. In the absence of intrinsic transitions ($\Omega=0$), the Hamiltonian is diagonal and the relaxation process reduces to pure decoherence. In this case, all eigenmodes and eigenvalues of the Lindbladian can be computed analytically (see Appendix B). When we apply our reset protocol with the reset state $|\delta\rangle=|0\rangle\langle 0|$, corresponding to the ground state to ensure $t_{s}(r)\geq 0$, the two slowest decaying modes form a complex-conjugate pair:
$c_{2}^{r}(t)=e^{-rt+i\phi}k,\quad c_{2}^{r,*}(t)=e^{-rt-i\phi}k.$ (14)
This overlap cannot be entirely eliminated in finite time, but it decays exponentially fast, leading to a relaxation timescale
$\tau_{{\rm rel}}\sim\frac{1}{r}+\frac{1}{|\text{Re}(\lambda_{3})|},$ (15)
where $1/r$ characterizes the reset protocol timescale, and $1/|\text{Re}(\lambda_{3})|$ represents the accelerated relaxation timescale with the slowest mode eliminated. Clearly, $\tau_{{\rm rel}}<1/|{\rm Re}(\lambda_{2})|$ for sufficiently large $r$. This structure is general and holds for more complex systems due to analogous spectral modifications introduced by the quantum reset. By appropriately tuning the reset rate, the relaxation timescale can approach $1/|\rm Re(\lambda_{3})|$ arbitrarily closely. Using Eq. (9), the full dynamics under our reset protocol can thus be determined.
Since the condition $c_{2}^{r}(t_{s})=0$ cannot be satisfied exactly in this case—though $|c_{2}^{r}(t_{s})|$ can be made exponentially small by choosing $t_{s}\sim 1/r$—we instead define $t_{s}$ such that a faster decaying mode is suppressed. This pragmatic approach effectively facilitates experimental implementation. Specifically, solving $c_{4}^{r}(t_{s})=0$ gives
$t_{s}=\frac{\ln\left[1+\frac{\Gamma_{1}k_{\text{tot}}}{r}\left(\frac{1}{\Gamma_{0}+\Gamma_{1}e^{(\beta_{0}-\beta_{\text{env}})E}}-\frac{1}{\Gamma_{0}+\Gamma_{1}}\right)\right]}{k_{\text{tot}}},$ (16)
where the total transition rate is $k_{\text{tot}}=r+\Gamma_{0}+\Gamma_{1}$. This choice eliminates the fastest decaying relaxation mode, ensuring the slowest mode becomes exponentially negligible for $t_{s}>1/r$. Importantly, alternative selections for $t_{s}$ can also achieve accelerated relaxation, indicating that detailed dynamics knowledge is not essential. Notably, if the initial inverse temperature satisfies $\beta_{0}<\beta_{\text{env}}$, a scenario analogous to “cooling”, the ground state $|0\rangle\langle 0|$ universally serves as an effective reset state regardless of the reset rate $r$. Conversely, when $\beta_{0}>\beta_{\text{env}}$, resetting to the ground state slows relaxation, necessitating instead the excited state $|1\rangle\langle 1|$ as the reset state for acceleration. This example demonstrates that reset-induced acceleration or deceleration of relaxation is readily controllable through appropriate state selection.
To illustrate the effect of the reset protocol, we next present numerical results. We characterize the distance between the transient state $\rho(t)$ and the stationary state $\rho_{ss}$ using the standard quantum trace distance
[…]
where $|A|:=\sqrt{A^{\dagger}A}$. Other measures, such as the L-infinity norm, yield qualitatively similar outcomes. Utilizing this metric, we illustrate the relaxation dynamics by plotting trace distances from $\rho(t)$ to equilibrium $\rho_{\text{eq}}$ over time $t$ for various rate $r$ in Figure 1 (a). The reset scenario employs initial conditions $\beta_{0}=2.0$, $k=0.32$ and $\phi=1$, whereas the non-reset scenario has $\beta_{0}=3.0,$ $k=0.21$ and $\phi=1$. With the environment at lower temperature $\beta_{\text{env}}=4.0$, the non-reset system starts closer to equilibrium. Nonetheless, strikingly, the reset protocol substantially accelerates relaxation, manifesting an anomalous effect reminiscent of the quantum Mpemba effect—systems initially farther from stationary state reach stationarity faster. We next plot the critical time $t_{s}$ required to close the reset as a function of the resetting rate $r$ in Figure 1 (b). The results demonstrate that $t_{s}$ decreases monotonically as $r$ increases, with all other parameters held constant. This finding indicates that reset protocols employing larger resetting rates facilitate more rapid relaxation of open quantum systems.
Figure 1: Relaxation dynamics of the two-state system. (a) and (c): Distances between $\rho(t)$ and $\rho_{\text{eq}}$ as the function of time $t$ with different resetting rate $r$. (b) and (d): The $t_{s}$ as a function of the resetting rate $r$. The parameters are chosen as $\beta_{0}=2.0$, $k=0.32$ and $\phi=1$ with reset protocol and $\beta_{0}=3.0,$ $k=0.21$ and $\phi=1$ without reset. For (a) and (b) $\Omega=0$, for (c) and (d) $\Omega=2.0$.
When intrinsic transitions exist ($\Omega\neq 0$), the reset protocol remains effective. In Figure 1 (c) confirms accelerated relaxation for $\Omega=2$, here measured by the L-infinity norm to reduce fluctuations in the curves.
Accelerating relaxation of quantum observables.—In experimental settings, observables rather than full density matrices are typically measured. Our protocol naturally extends to controlling the relaxation dynamics of arbitrary quantum observables. The time evolution of an observable $O$ under unperturbed dynamics reads:
$\langle O\rangle(t)={\rm Tr}[Oe^{\mathcal{L}t}(\rho_{0})]=\langle O\rangle_{ss}+\sum_{k=2}^{d^{2}}c_{k}e^{\lambda_{k}t}{\rm Tr}[OR_{k}].$ (17)
Applying our reset protocol, the modified observable dynamics becomes
$\langle O\rangle^{r}(t)$ $=\langle O\rangle_{ss}+\sum_{k}c_{k}^{r}(t)e^{\lambda_{k}t}\Theta(t_{s}-t){\rm Tr}[OR_{k}]$
$+\sum_{k}c_{k}^{r}(t_{s})e^{\lambda_{k}t}\Theta(t-t_{s}){\rm Tr}[OR_{k}].$ (18)
Thus, the same reset condition $c_{2}^{r}(t_{s})\approx 0$ also ensures fast convergence of $\langle O\rangle(t)$ to its steady-state value $\langle O\rangle_{ss}$. This allows for direct experimental application of our protocol at the level of observables, without requiring full state tomography.
Concluding remarks.—We have introduced a general framework for accelerating relaxation in open quantum systems via quantum reset dynamics, applicable to arbitrary initial states. The proposed protocol is simple to implement: it requires temporarily coupling the system to a reset channel and turning it off after a finite duration $t_{s}$, which can be tuned through the reset rate $r$. Additionally, the protocol can suppress multiple relaxation modes simultaneously, enabling enhanced control over the relaxation dynamics. In the future study, our approach could be extended to certain non-Markovian and periodically driven (Floquet) dissipative systems, and may offer a route to circumvent known trade-offs in relaxation control [18]. Overall, our results establish quantum reset as a versatile and experimentally feasible tool for controlling relaxation timescales in open quantum dynamics.
This work is partly supported by the Innovation Program for Quantum Science and Technology (2021ZD0303306), MOST(2022YFA1303100).
Appendix
Appendix A Derivation of the modified lindbladian
Acting the time-evolution operator generated by this generalized Lindblad operator on the eigenvalue-expanded form of the initial state
$\rho_{0}=\rho_{ss}+\sum_{k=2}^{d^{2}}\text{Tr}(L_{k}^{\dagger}\rho_{0})R_{k}$ (19)
results in
$\rho^{r}(t)$ $=e^{t\mathcal{L}_{r}}\rho_{0}$
$=e^{t\mathcal{L}_{r}}\rho_{ss}+\sum_{k=2}^{d^{2}}\text{Tr}(L_{k}^{\dagger}\rho_{0})e^{(\lambda_{k}-r)t}R_{k}.$ (20)
To get the detailed form of the first term, note that the stationary state under reset dynamics is given by
$\rho_{ss}^{r}=\lim_{t\rightarrow\infty}\rho^{r}(t)=\rho_{ss}+\sum_{k=2}^{d^{2}}\frac{r\langle\delta|L_{k}^{\dagger}|\delta\rangle}{r-\lambda_{k}}R_{k},$ (21)
and the stationary state has the property that
$e^{t\mathcal{L}_{r}}\rho_{ss}^{r}=\rho_{ss}^{r}.$ (22)
Thus, acting $e^{t\mathcal{L}_{r}}$ on both sides of the Eq.(21) and combining Eq.(22) bring about
$e^{t\mathcal{L}_{r}}\rho_{ss}+\sum_{k=2}^{d^{2}}\frac{r\langle\delta|L_{k}^{\dagger}|\delta\rangle}{r-\lambda_{k}}e^{(\lambda_{k}-r)t}R_{k}$
$=\rho_{ss}+\sum_{k=2}^{d^{2}}\frac{r\langle\delta|L_{k}^{\dagger}|\delta\rangle}{r-\lambda_{k}}R_{k}$ (23)
Appendix B Spectral analysis of the two-state model
The Eq. (13) can be rewritten as a matrix equation for the vector $\vec{\rho}(t)=[\rho_{00}(t),\rho_{01}(t),\rho_{10}(t),\rho_{11}(t)]^{T}$:
$\frac{d\vec{\rho}(t)}{dt}=L\vec{\rho}(t),$ (24)
where $L$ is a $4\times 4$ Liouvillian matrix in the Fock-Liouvillian space, taking the form as
$L=\left(\begin{array}[]{cccc}-\Gamma_{0}&-i\Omega&i\Omega&\Gamma_{1}\\ -i\Omega&-iE-\frac{\Gamma_{0}+\Gamma_{1}}{2}&0&i\Omega\\ i\Omega&0&iE-\frac{\Gamma_{0}+\Gamma_{1}}{2}&-i\Omega\\ \Gamma_{0}&i\Omega&-i\Omega&-\Gamma_{1}\end{array}\right).$ (25)
In the simple case $\Omega=0$, the right eigenvalues and eigenvectors ot the matrix $L$ are given by
$\lambda_{1}$ $=0,\ \vec{R}_{1}=(\Gamma_{1},0,0,\Gamma_{0})^{\text{T}}$
$\lambda_{4}$ $=-\Gamma_{0}-\Gamma_{1},\ \vec{R}_{2}=(-1,0,0,1)^{\text{T}}$
$\lambda_{2}$ $=\frac{1}{2}(-2i-\Gamma_{0}-\Gamma_{1}),\ \vec{R}_{3}=(0,1,0,0)^{\text{T}}$
$\lambda_{3}$ $=\frac{1}{2}(2i-\Gamma_{0}-\Gamma_{1}),\ \vec{R}_{4}=(0,0,1,0)^{\text{T}}$ (26)
Likewise, the left eigenvalues and eigenvectors are given by
$\lambda_{1}^{\star}$ $=0,\ \vec{L}_{1}=(1,0,0,1)^{\text{T}}$
$\lambda_{4}^{\star}$ $=-\Gamma_{0}-\Gamma_{1},\ \vec{L}_{2}=(-\Gamma_{0},0,0,\Gamma_{1})^{\text{T}}$
$\lambda_{2}^{\star}$ $=\frac{1}{2}(2i-\Gamma_{0}-\Gamma_{1}),\ \vec{L}_{3}=(0,1,0,0)^{\text{T}}$
$\lambda_{3}^{\star}$ $=\frac{1}{2}(-2i-\Gamma_{0}-\Gamma_{1}),\ \vec{L}_{4}=(0,0,1,0)^{\text{T}}.$ (27)
Then, by Eq.(8) in the main text, the expression for the critical time $t_{s}$ is given by
$t_{s}=\frac{\ln\left[1+\frac{\Gamma_{1}(r+\Gamma_{0}+\Gamma_{1})\left(\frac{1}{\Gamma_{0}+\Gamma_{0}e^{\beta_{0}E}}-\frac{1}{\Gamma_{0}+\Gamma_{1}}\right)}{r}\right]}{r+\Gamma_{0}+\Gamma_{1}},$ (28)
from which one can see that the critical time wouldn’t be affected by the coherence terms of the initial state.

### v3

@@
-Universal Acceleration of Quantum Relaxation by Reset-Induced Mpemba Effect
+Accelerating Quantum Relaxation via Temporary Reset: A Mpemba-Inspired Approach
 Abstract
 Slow relaxation processes spanning widely separated timescales pose fundamental challenges for probing steady-state properties and engineering functional quantum systems, such as quantum heat engines and quantum computing devices.
-We introduce a universal quantum reset protocol that enables exponential acceleration of relaxation in general Markovian open quantum systems.
-The protocol induces a quantum Mpemba effect, wherein systems initialized farther from equilibrium can relax faster than those closer to it, and crucially, this acceleration persists even when the slowest decaying Lindbladian modes form complex-conjugate pairs.
+We introduce a protocol that enables exponential acceleration of relaxation in general Markovian open quantum systems by temporarily coupling the system to a reset channel, inspired by the quantum Mpemba effect.
+Crucially, this acceleration persists even when the slowest decaying Lindbladian modes form complex-conjugate pairs.
 Unlike previous approaches, which typically target a single mode, our protocol eliminates multiple relaxation modes simultaneously.
-This framework provides a versatile and experimentally feasible tool for controlling quantum relaxation timescales, with broad implications for quantum thermodynamics, computation, and state preparation.
+Specifically, using a simple depolarizing channel, we show that the relaxation of a dissipative transverse-field Ising chain can be significantly accelerated, underscoring the experimental feasibility of the protocol.
+This framework provides a versatile tool for controlling quantum relaxation timescales, with broad implications for quantum thermodynamics, computation, and state preparation.
 Introduction.—Relaxation processes in open quantum systems, wherein a driven or interacting system settles into an equilibrium or a nonequilibrium steady state, are of fundamental importance in nonequilibrium physics and for practical quantum technologies.
@@
 This phenomenon implies the existence of an optimal initial state that minimizes the relaxation timescale.
-Theoretical frameworks for identifying such optimal states have been proposed in both classical [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31] and quantum systems [10, 32, 33, 34, 35, 36, 37, 38, 39], see also [40, 41] for recent reviews.
-However, existing quantum implementations typically impose unrealistic constraints, such as assuming pure initial states or requiring the second-largest eigenvalue of the Lindbladian to be real, which limit their applicability in general settings.
-Thus, a general and practical method for inducing faster relaxation toward arbitrary steady states in open quantum systems remains elusive.
-Likewise, a systematic recipe for prolonging the lifetime of metastable states, i.e., for slowing down the relaxation process is also lacking.
-In this study, we address these challenges by proposing an experimentally feasible protocol that can both accelerate and decelerate general quantum relaxation processes.
-The protocol is based on the quantum reset operation [42, 43, 44, 45, 46], which is a key component of many quantum computation tasks [47].
-For instance, qubit initialization to the computational ground state $|00...0\rangle$ is typically realized through a reset procedure.
-Recent experimental advancements have demonstrated the practical implementation of quantum reset, e.g.
-Aamir et al.
-have realized an autonomous quantum absorption refrigerator using superconducting circuits, capable of resetting a transmon qubit [48].
-Other reset schemes have also been proposed [49, 50, 51, 52].
+Theoretical frameworks for identifying such optimal states have been proposed in both classical [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32] and quantum systems [10, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42], see also [43, 44] for recent reviews.
+However, existing quantum implementations typically impose strict constraints, such as assuming pure initial states or requiring the second-largest eigenvalue of the Lindbladian to be real.
+Moreover, careful initial-state design is a delicate task, since it relies on detailed knowledge of the original initial state and the system dynamics, which is typically not known a prior in complex quantum systems.
+These issues limit their applicability in general settings.
+Thus, a general and practical method for inducing faster relaxation in open quantum systems, which is robust with respect to different initial states, remains elusive.
+In this study, we address these challenges by proposing an experimentally feasible protocol that can both accelerate and decelerate general quantum relaxation processes from any initial state, by resetting the system to a specified (possibly mixed) target state [45, 46, 47].
+Here, this type of quantum reset operation can be realized through engineered dissipation, i.e., quantum reservoir engineering [48, 49, 9, 50], and should be distinguished from the conventional reset typically used for qubit initialization in quantum computation [51, 52, 53, 54, 55, 56, 57, 58].
 In our protocol, we introduce a finite-duration reset phase into the open-system dynamics by temporarily coupling the system to a reset channel.
-We show that this operation can exponentially accelerate relaxation, even when starting from mixed initial states and in the presence of complex decay modes.
+We show that this protocol can exponentially accelerate relaxation, even when starting from mixed initial states and in the presence of complex decay modes.
 Remarkably, our method allows for the simultaneous elimination of multiple relaxation modes—a capability absent in previous approaches.
-Moreover, by applying the protocol selectively, the lifetime of metastable states can be extended, offering a systematic route to suppress relaxation.
+Moreover, selective application of the protocol enables the system of interest to remain in metastable states for extended periods, offering a systematic route to suppress relaxation.
 Setup.—We consider a general open quantum system defined on a Hilbert space $\mathcal{H}$ of dimension $d$, whose dynamics in the interaction picture are governed by the Lindblad-Gorini-Kossakowski-Sudarshan (LGKS) master equation $d\rho/dt=\mathcal{L}(\rho),$ where the Lindbladian superoperator $\mathcal{L}$ is given by
 $\mathcal{L}(\rho)\equiv-i[H,\rho]+\sum_{i}\left[J_{i}\rho J_{i}^{\dagger}-\frac{1}{2}\{J_{i}^{\dagger}J_{i},\rho\}\right].$ (1)
-[…]
+Here, $H$ is the Hamiltonian of the system and the jump operators $J_{i}$ describe the dissipative coupling to the environment.
+The evolution can be analyzed via the spectral decomposition of $\mathcal{L}$, whose eigenvalues are denoted $\lambda_{k}$ and ordered such that $0\geq\text{Re}(\lambda_{k})\geq\text{Re}(\lambda_{k+1})$.
+Assuming that $\mathcal{L}$ is diagonalizable, let $R_{k}$ and $L_{k}$ denote the corresponding right and left eigenmatrices, satisfying $\mathcal{L}(R_{k})=\lambda_{k}R_{k}$ and $\mathcal{L^{\dagger}}(L_{k})=\lambda_{k}^{\star}L_{k},\ k=1,...,d^{2}$, with biorthogonal normalization
 $\text{Tr}(L_{k}^{\dagger}R_{h})=\delta_{kh}.$ (2)
 The dual superoperator $\mathcal{L}^{\dagger}$ governs the Heisenberg-picture dynamics of observables:
-$\mathcal{L}^{\dagger}(O)=i[H,O]+\sum_{i}\left[J_{i}OJ_{i}^{\dagger}-\frac{1}{2}\{J_{i}^{\dagger}J_{i},O\}\right].$
+$\mathcal{L}^{\dagger}(O)=i[H,O]+\sum_{i}\left[J_{i}^{\dagger}OJ_{i}-\frac{1}{2}\{J_{i}^{\dagger}J_{i},O\}\right].$
 Given an initial state $\rho_{0}$, the system state at time $t$ evolves as
-$\rho(t)=e^{t\mathcal{L}}[\rho_{0}]=\rho_{ss}+\sum_{k=2}^{d^{2}}e^{\lambda_{k}t}c_{k}R_{k},$ (3)
+$\rho(t)=e^{t\mathcal{L}}[\rho_{0}]=\rho_{\rm ss}+\sum_{k=2}^{d^{2}}e^{\lambda_{k}t}c_{k}R_{k},$ (3)
 where $c_{k}\equiv\text{Tr}(L_{k}^{\dagger}\rho_{0})$.
-The unique stationary state of the open quantum system $\rho_{ss}$ is given by $\rho_{ss}=\lim_{t\rightarrow\infty}\rho(t)=R_{1}$, assuming $\lambda_{1}=0$ is non-degenerate.
-In the long-time limit, the relaxation is dominated by the slowest decaying mode, and the deviation from stationarity obeys
-$||\rho(t)-\rho_{ss}||\sim\exp\left(|\text{Re}\lambda_{2}|t\right),$
-so the characteristic relaxation timescale is given by $\tau_{{\rm{\rm rel}}}\sim\frac{1}{|\text{Re}(\lambda_{2})|}$.
+The unique stationary state of the open quantum system $\rho_{\rm ss}$ is given by $\rho_{\rm ss}=\lim_{t\rightarrow\infty}\rho(t)=R_{1}$, assuming $\lambda_{1}=0$ is non-degenerate.
+In the long-time limit, the relaxation is dominated by the slowest decaying mode, and the deviation from stationarity obeys $||\rho(t)-\rho_{\rm ss}||\sim\exp\left(-|\text{Re}\lambda_{2}|t\right),$ so the characteristic relaxation timescale is given by $\tau_{{\rm{\rm rel}}}\sim\frac{1}{|\text{Re}(\lambda_{2})|}$.
 For clarity of exposition, we first consider the simple case where the second eigenvalue $\lambda_{2}$ is real and non-degenerate, as in [10].
@@
 Reset protocol to accelerate relaxation processes.—Recognizing that the slowest decaying mode governs the relaxation timescale, we propose a quantum reset protocol that suppresses its excitation from a general initial state via a finite-duration reset phase, thereby enabling exponential acceleration of relaxation.
-Specifically, we let the system evolve under modified dynamics for a finite time $t_{s}$, during which it is stochastically reset to a chosen state $|\delta\rangle$ at random times, with events following a Poisson process of rate $r$.
-The reset state may also be taken as a mixed state $\rho_{\delta}$.
+Specifically, we let the system evolve under modified dynamics for a finite time $t_{s}$, during which it is stochastically reset to a chosen state $\rho_{\delta}$ at random times, with events following a Poisson process of rate $r$.
 After time $t_{s}$, the reset channel is turned off, and the system resumes evolving under the original Lindbladian $\mathcal{L}$.
-The modified dynamics during the reset phase is governed by a new Lindbladian $\mathcal{L}_{r}$, defined as
-$\mathcal{L}_{r}(\rho)=\mathcal{L}(\rho)+r\text{Tr}(\rho)|\delta\rangle\langle\delta|-r\rho.$ (4)
-This can be physically implemented by adding reset jump operators [42] $J_{i}^{r}=\sqrt{r}|\delta\rangle\langle\phi_{i}|$ to the Lindbladian (1), where $|\phi_{i}\rangle$ form an auxiliary complete orthonormal basis.
-When the reset state is a mixed state $\rho_{\delta}$, the new Lindbladian reads $\mathcal{L}_{r}(\rho)=\mathcal{L}(\rho)+r\text{Tr}(\rho)\rho_{\delta}-r\rho$.
-Importantly, this protocol modifies the relaxation spectrum without altering the eigenmodes.
-This is because the term $r\text{Tr}(\rho)|\delta\rangle\langle\delta|$ in the modified Lindbladian contributes only to the stationary mode and vanishes on all traceless eigenmodes $R_{i}$ with $i\geq 2$, which satisfy $\mathrm{Tr}(R_{i})=0$ due to $\mathrm{Tr}(L_{1}^{\dagger}R_{i})=0$ and $L_{1}^{\dagger}=\mathbb{I}$ [cf.
+That is, the system density matrix $\rho(t)$ evolves under the reset channel for $t\in[0,t_{s}]$ and decouples from the channel for $t>t_{s}$.
+The modified dynamics during the reset phase is governed by a new Lindbladian $\mathcal{L}_{r}$ [45], defined as
+$\mathcal{L}_{r}(\rho)=\mathcal{L}(\rho)+r\text{Tr}(\rho)\rho_{\delta}-r\rho.$ (4)
+Expressing $\rho_{\delta}=\sum_{\alpha}p_{\alpha}|\psi_{\alpha}\rangle\langle\psi_{\alpha}|$, $\mathcal{L}_{r}$ is equivalent to introducing jump operators $J_{i}^{r,\alpha}=\sqrt{rp_{\alpha}}|\psi_{\alpha}\rangle\langle\phi_{i}|$ into $\mathcal{L}$, where $|\phi_{i}\rangle$ form an auxiliary complete orthonormal basis (see End Matter).
+The term $r\text{Tr}(\rho)\rho_{\delta}$ contributes only to the stationary mode and vanishes on all traceless eigenmodes $R_{i}$ with $i\geq 2$, which satisfy $\mathrm{Tr}(R_{i})=0$ due to $\mathrm{Tr}(L_{1}^{\dagger}R_{i})=0$ and $L_{1}^{\dagger}=\mathbb{I}$ [cf.
 (2)].
-As a result, the modified Lindbladian acts as
-$\mathcal{L}_{r}(R_{i})=(\lambda_{i}-r)R_{i},\ i\in\{2,...,d^{2}\},$ (5)
-leading to a uniform spectral shift by $-r$ while preserving the mode structure.
+Consequently, the modified Lindbladian acts as
+$\mathcal{L}_{r}(R_{i})=(\lambda_{i}-r)R_{i},\ i\in\{2,...,d^{2}\}.$ (5)
+This induces a uniform spectral shift by $-r$ for all non-stationary eigenmodes, leaving the eigenmatrices ($R_{i}$, $i\geq 2$) themselves unchanged.
+In contrast, the stationary mode $R_{1}$ retains eigenvalue zero but is modified under reset.
 Let $\rho^{r}(t):=e^{t\mathcal{L}_{r}}\rho_{0}$ be the state at time $t$ under reset dynamics.
 Its spectral decomposition reads:
-$\rho^{r}(t)=$ $e^{t\mathcal{L}_{r}}\rho_{ss}+\sum_{k=2}^{d^{2}}\text{Tr}(L_{k}^{\dagger}\rho_{0})e^{(\lambda_{k}-r)t}R_{k}$
-::=:= $\rho_{ss}+\sum_{k}c_{k}^{r}(t)e^{\lambda_{k}t}R_{k},$ (6)
-where $c_{k}^{r}(t)\propto\text{Tr[$L_{k}^{\dagger}$$\rho^{r}(t)$]}$ are time-dependent mode amplitudes. Explicitly, these modified coefficients are given by
+$\rho^{r}(t)$ $=e^{t\mathcal{L}_{r}}\rho_{\rm ss}+\sum_{k=2}^{d^{2}}c_{k}e^{(\lambda_{k}-r)t}R_{k}$
+$:=\rho_{\rm ss}+\sum_{k}c_{k}^{r}(t)e^{\lambda_{k}t}R_{k},$ (6)
+where $c_{k}^{r}(t)$ are time-dependent mode amplitudes. Explicitly, these modified coefficients are given by (End Matter)
 $c_{k}^{r}(t)\equiv\left[c_{k}-\frac{r\cdot d_{k}}{r-\lambda_{k}}\right]e^{-rt}+\frac{r\cdot d_{k}}{r-\lambda_{k}}e^{-\lambda_{k}t},$ (7)
-where $d_{k}=\langle\delta|L_{k}^{\dagger}|\delta\rangle$.
-By properly choosing $t_{s}$, we can set $c_{2}^{r}(t_{s})=0$, thereby eliminating the overlap with the slowest decaying mode.
+where $d_{k}=\text{Tr}(L_{k}^{\dagger}\rho_{\delta})$.
+By properly choosing an optimal $t^{*}_{s}$, we can set $c_{2}^{r}(t^{*}_{s})=0$, thereby eliminating the overlap with the slowest decaying mode.
 Solving $c_{2}^{r}(t_{s})=0$ yields
-$t_{s}(r)=\frac{1}{r-\lambda_{2}}\ln\left[1-\frac{c_{2}(r-\lambda_{2})}{r\cdot d_{2}}\right].$ (8)
-This condition is always achievable by an appropriate choice of reset state $|\delta\rangle$ ensuring $t_{s}(r)\geq 0$.
+$t^{*}_{s}(r)=\frac{1}{r-\lambda_{2}}\ln\left[1-\frac{c_{2}(r-\lambda_{2})}{r\cdot d_{2}}\right].$ (8)
+By an appropriate choice of $\rho_{\delta}$ such that $c_{2}/d_{2}\leq 0$, $t^{*}_{s}(r)\geq 0$ can be assured.
 After the reset phase ($t>t_{s}$), the system returns to original Lindbladian evolution, now without the contribution from the slowest mode.
 The full evolution reads:
-$\rho^{r}(t)=$ $\rho_{ss}+\sum_{k}c_{k}^{r}(t)e^{\lambda_{k}t}\Theta(t_{s}-t)R_{k}$
+$\rho^{r}(t)=$ $\rho_{\rm ss}+\sum_{k}c_{k}^{r}(t)e^{\lambda_{k}t}\Theta(t_{s}-t)R_{k}$
 $+\sum_{k}c_{k}^{r}(t_{s})e^{\lambda_{k}t}\Theta(t-t_{s})R_{k},$ (9)
-where $\Theta(x)$ is the Heavside step function. This protocol reduces the relaxation time to
-$\tau_{{\rm rel}}\sim t_{s}+\frac{1}{|{\rm Re}(\lambda_{3})|},$ (10)
+where $\Theta(x)$ is the Heavside step function. This protocol with an optimal $t_{s}=t^{*}_{s}$ reduces the relaxation time to
+$\tau_{{\rm rel}}\sim t^{*}_{s}+\frac{1}{|{\rm Re}(\lambda_{3})|},$ (10)
 which is generically shorter than the unperturbed timescale $\tau_{{\rm{\rm rel}}}\sim\frac{1}{|\text{Re}(\lambda_{2})|}$.
+Eq.
+(9) can be generalized to the dynamics of observables [59], which may be more convenient to measure.
 Our protocol also enables controlled deceleration of relaxation.
-By choosing $|\delta\rangle$ and $t_{s}$ to amplify the magnitude of $c_{k}^{r}(t)$, one can extend the lifetime of selected metastable modes.
-This feature is especially useful in quantum metrology and computation, where long-lived metastable states play a functional role.
+By choosing $\rho_{\delta}$ and $t_{s}$ to amplify the magnitude of $c_{k}^{r}(t)$, one can prolong the time during which the system remains away from the steady state.
+This capability could be useful in quantum metrology and computation, where long-lived metastable states play a functional role.
 Our reset protocol thus provides a general framework for controlling relaxation processes in open quantum systems.
@@
 In this case, the long-time dynamics of the density matrix takes the form
-$\rho(t)=\rho_{ss}+e^{\lambda_{2}t}c_{2}R_{2}+e^{\lambda_{2}^{\dagger}t}c_{2}^{*}R_{2}^{\dagger}+...,$ (11)
+$\rho(t)=\rho_{\rm ss}+e^{\lambda_{2}t}c_{2}R_{2}+e^{\lambda_{2}^{\dagger}t}c_{2}^{*}R_{2}^{\dagger}+\mathcal{O}(e^{\lambda_{3}t}),$ (11)
 where $c_{2}=\text{Tr}(L_{2}^{\dagger}\rho_{0})$ and $c_{2}^{*}=\text{Tr}(L_{2}\rho_{0})$ is its complex conjugate.
@@
 Our reset protocol, however, naturally addresses this limitation. The modified coefficients under quantum reset are given by
-$c_{2}^{r}(t)=\left[c_{2}-\frac{r\cdot d_{2}}{r-\lambda_{2}}\right]e^{-rt}+\frac{r\cdot d_{2}}{r-\lambda_{2}}e^{-\lambda_{2}t}$ ,
-$c_{2}^{r,*}(t)=\left[c_{2}^{*}-\frac{r\cdot d_{2}^{*}}{r-\lambda_{2}^{*}}\right]e^{-rt}+\frac{r\cdot d_{2}^{*}}{r-\lambda_{2}^{*}}e^{-\lambda_{2}^{*}t}$
+$c_{2}^{r}(t)=\left[c_{2}-\frac{r\cdot d_{2}}{r-\lambda_{2}}\right]e^{-rt}+\frac{r\cdot d_{2}}{r-\lambda_{2}}e^{-\lambda_{2}t}$ , (12a)
+$c_{2}^{r,*}(t)=\left[c_{2}^{*}-\frac{r\cdot d_{2}^{*}}{r-\lambda_{2}^{*}}\right]e^{-rt}+\frac{r\cdot d_{2}^{*}}{r-\lambda_{2}^{*}}e^{-\lambda_{2}^{*}t}$ (12b)
 Importantly, we observe that these coefficients remain complex conjugates of each other for all $t$, i.e., $[c_{2}^{r}(t)]^{*}=c_{2}^{r,*}(t)$.
@@
 because $c_{2}^{r}(t_{s})=0\Rightarrow c_{2}^{r,*}(t_{s})=[c_{2}^{r}(t_{s})]^{*}=0$.
-This condition can be fulfilled by tuning the reset rate $r$ and the reset duration $t_{s}$, thereby simultaneously eliminating both components of the dominant complex relaxation mode.
-In practice, it suffices to reduce $|c_{2}^{r}(t_{s})|$ to an exponentially small value ($\text{Re}[c_{2}^{r}(t_{s})]\ll 1,\ \text{Im}[c_{2}^{r}(t_{s})]\ll 1$), which still leads to exponential acceleration of relaxation.
-We demonstrate this behavior concretely in the following example.
-Example.—To illustrate our protocol, we consider a minimal example: a two-level quantum system, which may represent a qubit or a spin-$1/2$ particle.
-We set the ground state energy $E_{0}=0$ and excited state energy $E_{1}=E$, such that the system Hamiltonian reads
-$H=E|1\rangle\langle 1|+\Omega(|0\rangle\langle 1|+|1\rangle\langle 0|)\text{,}$
-where $\Omega$ is the intrinsic coherent coupling between the two levels.
-The system is coupled to a thermal environment through the Lindblad jump operators $J_{0}=|1\rangle\langle 0|=\sigma_{+}$ and $J_{1}=|0\rangle\langle 1|=\sigma_{-}$.
-The resulting open-system dynamics are governed by the master equation:
-$\frac{d\rho(t)}{dt}=$ $-i[H,\rho]+\Gamma_{0}\left(J_{0}\rho J_{1}-\frac{1}{2}\{J_{0}J_{1},\rho\}\right)$
-$+\Gamma_{1}\left(J_{1}\rho J_{0}-\frac{1}{2}\{J_{1}J_{0},\rho\}\right),$ (13)
-where $\Gamma_{0,1}$ are the transition rates from state $|0\rangle$ to state $|1\rangle$ or converse.
-Assuming the rates satisfy detailed balance with inverse temperature $\Gamma_{0}e^{-\beta_{\text{env}}E_{0}}=\Gamma_{1}e^{-\beta_{\text{env}}E_{1}}$, i.e., $\Gamma_{0}=\Gamma_{1}e^{-\beta_{\text{env}}E},$ the system evolves to a unique equilibrium state $\rho_{eq}=\frac{e^{-\beta_{\text{env}}H}}{\text{Tr}(e^{-\beta_{\text{env}}H})}$ irrespective of the initial condition.
+In some cases, this condition can be exactly satisfied by tuning $r$ and $t_{s}$.
+Although Eq.
+(13) does not always admit an exact solution, in practice, it suffices to reduce $|c_{2}^{r}(t_{s})|$ to an exponentially small value ($\text{Re}[c_{2}^{r}(t_{s})]\ll 1,\ \text{Im}[c_{2}^{r}(t_{s})]\ll 1$), which still leads to exponential acceleration of relaxation.
+Additionally, we show in [59] that, if the condition
+${\rm Re}(c_{2}^{*}d_{2})\leq|c_{2}|^{2}$ (14)
+is satisfied, $|c_{2}^{r}(t_{s})|<|c_{2}|$ can always be achieved for any positive $t_{s}$ that below a certain threshold.
+This is a weak condition that be readily satisfied for many choices of $\rho_{\delta}$.
+It is strictly weaker than the commonly satisfied condition $c_{2}/d_{2}\leq 0$, which, in the case where $\lambda_{2}$ is real, guarantees the existence of a $t_{s}^{*}$ such that $c^{r}_{2}(t_{s}^{*})=0$.
+Eq.
+(14) thus provides a sufficient condition for reset-induced Mpemba-type acceleration [20].
+It implies that fine-tuning of either $\rho_{\delta}$ or $t_{s}$ is unnecessary for achieving acceleration.
+We illustrate this behavior concretely in the following examples.
+Example 1.—To illustrate our protocol, we first consider a minimal example: a two-level quantum system, which may represent a qubit or a spin-$1/2$ particle.
+While minimal, this example is crucial in many practical settings, as qubit reset often constitutes the bottleneck in large-scale quantum processes [60, 58].
+We set the ground state energy $E_{0}=0$ and excited state energy $E_{1}=E$, such that the system Hamiltonian reads $H=E|1\rangle\langle 1|+\Omega(\sigma^{+}+\sigma^{-})$, where $\Omega$ is the intrinsic coherent coupling between the two levels, $\sigma^{+}=|1\rangle\langle 0|$ and $\sigma^{-}=|0\rangle\langle 1|$.
+The system is coupled to a thermal environment through the Lindblad jump operators $J_{0}=\sqrt{\gamma_{0}}\sigma^{+}$ and $J_{1}=\sqrt{\gamma_{1}}\sigma_{-}$.
+Here, $\gamma_{0}$ ($\gamma_{1}$) is the transition rate from state $|0\rangle$ to state $|1\rangle$ (or converse).
+Assuming the rates satisfy detailed balance with inverse temperature $\gamma_{0}e^{-\beta_{\text{env}}E_{0}}=\gamma_{1}e^{-\beta_{\text{env}}E_{1}}$, i.e., $\gamma_{0}=\gamma_{1}e^{-\beta_{\text{env}}E},$ the system evolves to a unique equilibrium state $\rho_{\rm eq}=\frac{e^{-\beta_{\text{env}}H}}{\text{Tr}(e^{-\beta_{\text{env}}H})}$ irrespective of the initial condition.
 We initialize the system in a general mixed state:
@@
 In the absence of intrinsic transitions ($\Omega=0$), the Hamiltonian is diagonal and the relaxation process reduces to pure decoherence.
-In this case, all eigenmodes and eigenvalues of the Lindbladian can be computed analytically (see Appendix B).
-When we apply our reset protocol with the reset state $|\delta\rangle=|0\rangle\langle 0|$, corresponding to the ground state to ensure $t_{s}(r)\geq 0$, the two slowest decaying modes form a complex-conjugate pair:
+In this case, all eigenmodes and eigenvalues of the Lindbladian can be computed analytically [59].
+When we apply our reset protocol with the reset state $\rho_{\delta}=|0\rangle\langle 0|$, corresponding to the ground state to ensure $t_{s}(r)\geq 0$, the two slowest decaying modes form a complex-conjugate pair:
 $c_{2}^{r}(t)=e^{-rt+i\phi}k,\quad c_{2}^{r,*}(t)=e^{-rt-i\phi}k.$ (15)
@@
 Clearly, $\tau_{{\rm rel}}<1/|{\rm Re}(\lambda_{2})|$ for sufficiently large $r$.
-This structure is general and holds for more complex systems due to analogous spectral modifications introduced by the quantum reset.
 By appropriately tuning the reset rate, the relaxation timescale can approach $1/|\rm Re(\lambda_{3})|$ arbitrarily closely.
@@
 Since the condition $c_{2}^{r}(t_{s})=0$ cannot be satisfied exactly in this case—though $|c_{2}^{r}(t_{s})|$ can be made exponentially small by choosing $t_{s}\sim 1/r$—we instead define $t_{s}$ such that a faster decaying mode is suppressed.
-This pragmatic approach effectively facilitates experimental implementation.
 Specifically, solving $c_{4}^{r}(t_{s})=0$ gives
@@
 where the total transition rate is $k_{\text{tot}}=r+\gamma_{0}+\gamma_{1}$.
-This choice eliminates the fastest decaying relaxation mode, ensuring the slowest mode becomes exponentially negligible for $t_{s}>1/r$.
-Importantly, alternative selections for $t_{s}$ can also achieve accelerated relaxation, indicating that detailed dynamics knowledge is not essential.
-Notably, if the initial inverse temperature satisfies $\beta_{0}<\beta_{\text{env}}$, a scenario analogous to “cooling”, the ground state $|0\rangle\langle 0|$ universally serves as an effective reset state regardless of the reset rate $r$.
-Conversely, when $\beta_{0}>\beta_{\text{env}}$, resetting to the ground state slows relaxation, necessitating instead the excited state $|1\rangle\langle 1|$ as the reset state for acceleration.
-This example demonstrates that reset-induced acceleration or deceleration of relaxation is readily controllable through appropriate state selection.
-To illustrate the effect of the reset protocol, we next present numerical results.
-We characterize the distance between the transient state $\rho(t)$ and the stationary state $\rho_{ss}$ using the standard quantum trace distance
-[…]
-where $|A|:=\sqrt{A^{\dagger}A}$.
-Other measures, such as the L-infinity norm, yield qualitatively similar outcomes.
-Utilizing this metric, we illustrate the relaxation dynamics by plotting trace distances from $\rho(t)$ to equilibrium $\rho_{\text{eq}}$ over time $t$ for various rate $r$ in Figure 1 (a).
-The reset scenario employs initial conditions $\beta_{0}=2.0$, $k=0.32$ and $\phi=1$, whereas the non-reset scenario has $\beta_{0}=3.0,$ $k=0.21$ and $\phi=1$.
-With the environment at lower temperature $\beta_{\text{env}}=4.0$, the non-reset system starts closer to equilibrium.
-Nonetheless, strikingly, the reset protocol substantially accelerates relaxation, manifesting an anomalous effect reminiscent of the quantum Mpemba effect—systems initially farther from stationary state reach stationarity faster.
-We next plot the critical time $t_{s}$ required to close the reset as a function of the resetting rate $r$ in Figure 1 (b).
-The results demonstrate that $t_{s}$ decreases monotonically as $r$ increases, with all other parameters held constant.
-This finding indicates that reset protocols employing larger resetting rates facilitate more rapid relaxation of open quantum systems.
+This choice fully eliminates the fastest decaying relaxation mode.
+In addition, for a sufficiently large reset rate $r$, the slowest mode’s amplitude becomes exponentially small after the reset phase.
+Alternative selections for $t_{s}$ can also achieve accelerated relaxation, indicating that detailed dynamics knowledge is not essential.
+Notably, if the initial inverse temperature satisfies $\beta_{0}<\beta_{\text{env}}$, a scenario analogous to “cooling”, the ground state $|0\rangle\langle 0|$ serves as a useful $\rho_{\delta}$ regardless of the rate $r$.
+Conversely, when $\beta_{0}>\beta_{\text{env}}$, resetting to the ground state decelerates relaxation, necessitating the choice $\rho_{\delta}=|1\rangle\langle 1|$ for acceleration.
+This example demonstrates that, by appropriately choosing $\rho_{\delta}$, one can switch between reset-induced acceleration and deceleration of relaxation.
+We next present numerical results for illustration.
+We characterize the distance between the transient state $\rho(t)$ and the stationary state $\rho_{\rm ss}$ using the standard trace distance $D[\rho(t)|\rho_{\rm ss}]:=\text{Tr}|\rho(t)-\rho_{\rm ss}|/2$ where $|A|:=\sqrt{A^{\dagger}A}$.
+Other measures, such as the $L_{\infty}$ norm $D_{\infty}[\rho(t)|\rho_{\rm ss}]:=\max_{i}|\lambda_{i}|$ [$\lambda_{i}$ is the $i$-th eigenvalue of $\rho(t)-\rho_{\rm ss}$], yield qualitatively similar outcomes.
+We first plot $D[\rho(t)|\rho_{\rm eq}]$ over $t$ for various rate $r$ in Fig.
+1 (a).
+The parameters for $r>0$ cases are: initial conditions $\beta_{0}=2.0$, $k=0.32$ and $\phi=1$; The $r=0$ scenario has $\beta_{0}=3.0,$ $k=0.21$ and $\phi=1$.
+The environment is at lower temperature $\beta_{\text{env}}=4.0$.
+Although the $r>0$ cases are initially farther from equilibrium compared to the $r=0$ case, the reset protocol substantially accelerates relaxation so that the former cases reach equilibrium even faster.
+We next plot $t_{s}$ as a function of $r$ in Fig.
+1 (b) and (d), which show that $t_{s}$ decreases monotonically as $r$ increases.
+Fig.
+1 (c) confirms accelerated relaxation for $\Omega=2$, here measured by the $L_{\infty}$ norm to reduce fluctuations in the curves.
 Figure 1: Relaxation dynamics of the two-state system.
@@
 For (a) and (b) $\Omega=0$, for (c) and (d) $\Omega=2.0$.
-When intrinsic transitions exist ($\Omega\neq 0$), the reset protocol remains effective.
-In Figure 1 (c) confirms accelerated relaxation for $\Omega=2$, here measured by the L-infinity norm to reduce fluctuations in the curves.
-Accelerating relaxation of quantum observables.—In experimental settings, observables rather than full density matrices are typically measured.
-Our protocol naturally extends to controlling the relaxation dynamics of arbitrary quantum observables.
-The time evolution of an observable $O$ under unperturbed dynamics reads:
-$\langle O\rangle(t)={\rm Tr}[Oe^{\mathcal{L}t}(\rho_{0})]=\langle O\rangle_{ss}+\sum_{k=2}^{d^{2}}c_{k}e^{\lambda_{k}t}{\rm Tr}[OR_{k}].$ (17)
-Applying our reset protocol, the modified observable dynamics becomes
-$\langle O\rangle^{r}(t)$ $=\langle O\rangle_{ss}+\sum_{k}c_{k}^{r}(t)e^{\lambda_{k}t}\Theta(t_{s}-t){\rm Tr}[OR_{k}]$
-$+\sum_{k}c_{k}^{r}(t_{s})e^{\lambda_{k}t}\Theta(t-t_{s}){\rm Tr}[OR_{k}].$ (18)
-Thus, the same reset condition $c_{2}^{r}(t_{s})\approx 0$ also ensures fast convergence of $\langle O\rangle(t)$ to its steady-state value $\langle O\rangle_{ss}$.
-This allows for direct experimental application of our protocol at the level of observables, without requiring full state tomography.
-Concluding remarks.—We have introduced a general framework for accelerating relaxation in open quantum systems via quantum reset dynamics, applicable to arbitrary initial states.
-The proposed protocol is simple to implement: it requires temporarily coupling the system to a reset channel and turning it off after a finite duration $t_{s}$, which can be tuned through the reset rate $r$.
-Additionally, the protocol can suppress multiple relaxation modes simultaneously, enabling enhanced control over the relaxation dynamics.
-In the future study, our approach could be extended to certain non-Markovian and periodically driven (Floquet) dissipative systems, and may offer a route to circumvent known trade-offs in relaxation control [18].
-Overall, our results establish quantum reset as a versatile and experimentally feasible tool for controlling relaxation timescales in open quantum dynamics.
+Example 2.—To illustrate the potential feasibility of our protocol in complex quantum systems, we consider another example, the dissipative transverse-field Ising model (TFIM), a paradigmatic system relevant to various quantum platforms.
+We numerically study an open dissipative TFIM of length $N$ with Hamiltonian
+$H=-J\sum_{i=1}^{N-1}\sigma_{i}^{z}\sigma_{i+1}^{z}-g\sum_{i=1}^{N}\sigma_{i}^{x},$ (18)
+and jump operators $J_{i,\downarrow}=\sqrt{\gamma}\,\sigma_{i}^{-},\quad J_{i,\uparrow}=\sqrt{\gamma e^{-\beta}}\;\sigma_{i}^{+}$.
+Here, $\sigma^{x,z}_{i}$ are Pauli operators, $\sigma_{i}^{-}=|0\rangle\langle 1|$ and $\sigma_{i}^{+}=|1\rangle\langle 0|$.
+We fix $J=1$, $g=1.2$, $\gamma=0.5$, $\beta=1/k_{B}T=1$ throughout.
+We choose $\rho_{\delta}=\mathbb{I}/d$ (with $d=2^{N}$).
+The initial state $\rho_{0}$ is chosen by normalizing $\rho_{\rm ss}+\alpha V_{2}$, with $\rho_{\rm ss}$ being the steady-state and $V_{2}$ being the second normalized eigenvector of the Lindbladian.
+A larger $\alpha$ implies that $\rho_{0}$ is farther from $\rho_{\rm ss}$.
+Here, $\rho_{\rm ss}$ is nontrivial, with nonzero entanglement.
+Parameters: $t_{s}=0.5\,\tau_{2}$ and $t_{s}=0.08\,\tau_{2}$ with $\tau_{2}=1/|\operatorname{Re}\lambda_{2}|$, the relaxation timescale, total duration $T=6\,\tau_{2}$, $\alpha=0.55/0.05$.
+Notably, $t_{s}$ is chosen arbitrarily, rather than being determined a prior.
+The accelerated relaxation processes under different reset rate are shown in Fig.
+2.
+We also examine cases with different parameters in [59], where acceleration is consistently observed.
+In Fig.
+2 (b) and (d), genuine Mpemba effects are observed, where the system is first driven away from the steady state by the reset protocol compared to the no-reset case, and then reaches stationarity faster.
+The existence of such an effect broadens the range of usable reset states.
+Some remarks are in order regarding experimental implications.
+In general, the optimal duration $t_{s}$ of the reset phase cannot be determined a priori from theoretical considerations, as the relevant dynamical details are typically unknown in complex quantum systems.
+The TFIM example shows how this challenge can be overcome: fine-tuning of $t_{s}$ is not required—any finite value shorter than the native relaxation time leads to substantial speed-up, and $t_{s}$ can be further adjusted a posteriori based on observed dynamics.
+This robustness, without requiring prior knowledge of the system dynamics and initial states, distinguishes our protocol from previous approaches.
+Additionally, the reset channel in the example corresponds to a depolarizing channel, which is conceptually straightforward to implement, e.g., by applying an isotropic white-noise field or coupling an infinite-temperature reservoir.
+Figure 2: Acceleration of the relaxation of TFIMs.
+Reset rate $r$ are chosen as $0,1.0,5.0,10.0,20.0$.
+(a) $N=5$ ($d=32$), $\alpha=0.55$, $t_{s}=0.08\tau_{2}$.
+(b) $N=5$, $\alpha=0.05$, $t_{s}=0.50\tau_{2}$.
+(c) $N=6$ ($d=64$), $\alpha=0.55$, $t_{s}=0.08\tau_{2}$.
+(d) $N=6$, $\alpha=0.05$, $t_{s}=0.50\tau_{2}$.
+The vertical gray lines indicate $t_{s}$ when the reset channel is removed.
+Inset: Schematic representation of the dissipative TFIM, with each spin coupled to a thermal bath.
+Experimental feasibility in general cases.—The TFIM example illustrates the power of our protocol: the acceleration of relaxation via a depolarizing channel has the potential to be applied to complex quantum platforms.
+However, in general, it may be necessary to choose other $\rho_{\delta}$ to achieve acceleration.
+For few-body systems (such as single Rydberg atoms or other multi-level emitters), the protocol may be implemented with quantum reservoir engineering.
+Such techniques have become increasingly routine for tailoring dissipative dynamics in diverse quantum systems, including atoms [61, 62, 63], trapped ions [48, 64], superconducting circuits [61, 50, 65], and optomechanical setups [66, 67, 68].
+For quantum many-body systems, an exact implementation via reservoir engineering generally requires high-order interactions.
+Such engineered $N$-body dissipators have been put forward theoretically [69, 70] and realized experimentally up to four-body interactions [71].
+A complementary route is to approximate the reset channel: combine any available state-preparation protocol for $\rho_{\delta}$, whether reservoir engineering, measurement-based feedback, or other methods, with Trotterization [72, 73, 74] (see End Matter).
+This strategy sidesteps the need to engineer reset jump operators directly and therefore avoids high-order couplings even in many-body systems.
+The main experimental challenge then reduces to preparing $\rho_{\delta}$.
+Encouragingly, fast and high-fidelity state preparation has been reported across diverse platforms [75, 76, 63, 77].
+Notably, Ref.
+[77] realized state preparation in a 35-spin TFIM with Trotterization.
+This suggests that our protocol may be testable and applicable on existing platforms.
+Concluding remarks.—We have introduced a general framework for accelerating relaxation in open quantum systems via temporary reset, applicable to arbitrary initial states.
+The proposed protocol can suppress multiple relaxation modes simultaneously, enabling enhanced control over the relaxation dynamics.
+As a practical example, we demonstrated that the relaxation of a TFIM can be significantly accelerated by using a simple depolarizing channel.
+This example highlights a powerful feature of our protocol: leveraging easily prepared states to accelerate the preparation of states that are difficult to reach.
+Furthermore, introducing temporary dephasing noise—a different type of channel—can also accelerate relaxation (End Matter).
+This suggests the broader applicability of our central idea: temporarily coupling the system to various quantum channels may provide a general route to enhanced relaxation.
+In future studies, our approach could be extended to certain non-Markovian and Floquet dissipative systems (e.g., using ideas in [10, 78]).
+Overall, our results establish temporary reset as a powerful and experimentally feasible tool for controlling relaxation timescales in open quantum dynamics.
 This work is partly supported by the Innovation Program for Quantum Science and Technology (2021ZD0303306), MOST(2022YFA1303100).
-Appendix
-Appendix A Derivation of the modified lindbladian
-Acting the time-evolution operator generated by this generalized Lindblad operator on the eigenvalue-expanded form of the initial state
-$\rho_{0}=\rho_{ss}+\sum_{k=2}^{d^{2}}\text{Tr}(L_{k}^{\dagger}\rho_{0})R_{k}$ (19)
-results in
-$\rho^{r}(t)$ $=e^{t\mathcal{L}_{r}}\rho_{0}$
-$=e^{t\mathcal{L}_{r}}\rho_{ss}+\sum_{k=2}^{d^{2}}\text{Tr}(L_{k}^{\dagger}\rho_{0})e^{(\lambda_{k}-r)t}R_{k}.$ (20)
-To get the detailed form of the first term, note that the stationary state under reset dynamics is given by
-$\rho_{ss}^{r}=\lim_{t\rightarrow\infty}\rho^{r}(t)=\rho_{ss}+\sum_{k=2}^{d^{2}}\frac{r\langle\delta|L_{k}^{\dagger}|\delta\rangle}{r-\lambda_{k}}R_{k},$ (21)
-and the stationary state has the property that
-$e^{t\mathcal{L}_{r}}\rho_{ss}^{r}=\rho_{ss}^{r}.$ (22)
-Thus, acting $e^{t\mathcal{L}_{r}}$ on both sides of the Eq.(21) and combining Eq.(22) bring about
-$e^{t\mathcal{L}_{r}}\rho_{ss}+\sum_{k=2}^{d^{2}}\frac{r\langle\delta|L_{k}^{\dagger}|\delta\rangle}{r-\lambda_{k}}e^{(\lambda_{k}-r)t}R_{k}$
-$=\rho_{ss}+\sum_{k=2}^{d^{2}}\frac{r\langle\delta|L_{k}^{\dagger}|\delta\rangle}{r-\lambda_{k}}R_{k}$ (23)
-Appendix B Spectral analysis of the two-state model
-The Eq. (13) can be rewritten as a matrix equation for the vector $\vec{\rho}(t)=[\rho_{00}(t),\rho_{01}(t),\rho_{10}(t),\rho_{11}(t)]^{T}$:
-$\frac{d\vec{\rho}(t)}{dt}=L\vec{\rho}(t),$ (24)
-where $L$ is a $4\times 4$ Liouvillian matrix in the Fock-Liouvillian space, taking the form as
-$L=\left(\begin{array}[]{cccc}-\Gamma_{0}&-i\Omega&i\Omega&\Gamma_{1}\\ -i\Omega&-iE-\frac{\Gamma_{0}+\Gamma_{1}}{2}&0&i\Omega\\ i\Omega&0&iE-\frac{\Gamma_{0}+\Gamma_{1}}{2}&-i\Omega\\ \Gamma_{0}&i\Omega&-i\Omega&-\Gamma_{1}\end{array}\right).$ (25)
-In the simple case $\Omega=0$, the right eigenvalues and eigenvectors ot the matrix $L$ are given by
-$\lambda_{1}$ $=0,\ \vec{R}_{1}=(\Gamma_{1},0,0,\Gamma_{0})^{\text{T}}$
-$\lambda_{4}$ $=-\Gamma_{0}-\Gamma_{1},\ \vec{R}_{2}=(-1,0,0,1)^{\text{T}}$
-$\lambda_{2}$ $=\frac{1}{2}(-2i-\Gamma_{0}-\Gamma_{1}),\ \vec{R}_{3}=(0,1,0,0)^{\text{T}}$
-$\lambda_{3}$ $=\frac{1}{2}(2i-\Gamma_{0}-\Gamma_{1}),\ \vec{R}_{4}=(0,0,1,0)^{\text{T}}$ (26)
-Likewise, the left eigenvalues and eigenvectors are given by
-$\lambda_{1}^{\star}$ $=0,\ \vec{L}_{1}=(1,0,0,1)^{\text{T}}$
-$\lambda_{4}^{\star}$ $=-\Gamma_{0}-\Gamma_{1},\ \vec{L}_{2}=(-\Gamma_{0},0,0,\Gamma_{1})^{\text{T}}$
-$\lambda_{2}^{\star}$ $=\frac{1}{2}(2i-\Gamma_{0}-\Gamma_{1}),\ \vec{L}_{3}=(0,1,0,0)^{\text{T}}$
-$\lambda_{3}^{\star}$ $=\frac{1}{2}(-2i-\Gamma_{0}-\Gamma_{1}),\ \vec{L}_{4}=(0,0,1,0)^{\text{T}}.$ (27)
-Then, by Eq.(8) in the main text, the expression for the critical time $t_{s}$ is given by
-$t_{s}=\frac{\ln\left[1+\frac{\Gamma_{1}(r+\Gamma_{0}+\Gamma_{1})\left(\frac{1}{\Gamma_{0}+\Gamma_{0}e^{\beta_{0}E}}-\frac{1}{\Gamma_{0}+\Gamma_{1}}\right)}{r}\right]}{r+\Gamma_{0}+\Gamma_{1}},$ (28)
-from which one can see that the critical time wouldn’t be affected by the coherence terms of the initial state.
+End Matter
+Derivation of the modified Lindbladian $\mathcal{L}_{r}$.—Here, we derive Eq.
+(4) of the main text.
+Let $\{|\phi_{i}\rangle\}_{i=1}^{d}$ be an orthonormal basis, satisfying $\sum_{i}|\phi_{i}\rangle\!\langle\phi_{i}|=\mathbb{I}$.
+The reset state is generally written as $\rho_{\delta}=\sum_{\alpha}p_{\alpha}|\psi_{\alpha}\rangle\!\langle\psi_{\alpha}|$.
+Introduce jump operators
+$J_{i,\alpha}^{r}=\sqrt{r\,p_{\alpha}}|\psi_{\alpha}\rangle\!\langle\phi_{i}|,\quad i=1,\dots,d.$
+The associated Lindblad term is
+$\mathcal{D}_{r}(\rho)=\sum_{i=1}^{d}\!\bigl{(}J_{i}^{r}\rho\,J_{i}^{r\dagger}-\tfrac{1}{2}\{J_{i}^{r\dagger}J_{i}^{r},\rho\}\bigr{)}.$ (19)
+Direct evaluation gives
+$\sum_{i,\alpha}J_{i,\alpha}^{r}\rho\,J_{i,\alpha}^{r\dagger}=r\,\rho_{\delta}\,\text{Tr}[\rho]$
+$-\tfrac{1}{2}\sum_{i,\alpha}\{J_{i,\alpha}^{r\dagger}J_{i,\alpha}^{r},\rho\}=-r\,\rho,$
+where $\text{Tr}[\rho]=\sum_{i}\langle\phi_{i}|\rho|\phi_{i}\rangle$, $\sum_{i}|\phi_{i}\rangle\!\langle\phi_{i}|=\mathbb{I}$ and $\langle\psi_{\alpha}|\psi_{\alpha}\rangle=1$ have been used.
+This yields $\mathcal{D}_{r}(\rho)=r\bigl{[}\text{Tr}[\rho]\,\rho_{\delta}-\rho\bigr{]}$.
+Hence, we arrive at Eq.
+(4),
+$\mathcal{L}_{r}(\rho)=\mathcal{L}(\rho)+r\text{Tr}[\rho]\,\rho_{\delta}-r\rho.$ (20)
+Derivation of the modified coefficients.—Here, we derive Eq. (7), the explicit form of the modified coefficients with reset. Acting the modified semi-group $e^{t\mathcal{L}_{r}}$ on both sides of
+$\rho_{0}=\rho_{\rm ss}+\sum_{k=2}^{d^{2}}\text{Tr}(L_{k}^{\dagger}\rho_{0})R_{k}$ (21)
+yields
+$\rho^{r}(t)=e^{t\mathcal{L}_{r}}\rho_{0}=e^{t\mathcal{L}_{r}}\rho_{\rm ss}+\sum_{k=2}^{d^{2}}\text{Tr}(L_{k}^{\dagger}\rho_{0})e^{(\lambda_{k}-r)t}R_{k}.$ (22)
+To proceed, we use an explicit form of the steady state under reset dynamics [45]
+$\rho_{\rm ss}^{r}=\lim_{t\rightarrow\infty}\rho^{r}(t)=\rho_{\rm ss}+\sum_{k=2}^{d^{2}}\frac{rd_{k}}{r-\lambda_{k}}R_{k}.$ (23)
+Then, expressing the steady-state condition $e^{t\mathcal{L}_{r}}\rho_{\rm ss}^{r}=\rho_{\rm ss}^{r}$ with Eq. (23) and applying Eq. (5), we get
+$e^{t\mathcal{L}_{r}}\rho_{\rm ss}+\sum_{k=2}^{d^{2}}\frac{rd_{k}}{r-\lambda_{k}}e^{(\lambda_{k}-r)t}R_{k}=\rho_{\rm ss}+\sum_{k=2}^{d^{2}}\frac{rd_{k}}{r-\lambda_{k}}R_{k},$
+from which we solve $e^{t\mathcal{L}_{r}}\rho_{\rm ss}$.
+Substituting $e^{t\mathcal{L}_{r}}\rho_{\rm ss}$ into Eq.
+(22) and comparing with Eq.
+(Accelerating Quantum Relaxation via Temporary Reset: A Mpemba-Inspired Approach), we identify modified coefficients $c_{k}^{r}(t)$ defined in Eq.
+(7).
+Realizing the reset protocol with arbitrary state preparation processes.—We now show that any state preparation protocol could approximately realize the reset channel in our scheme via a Trotterization-based construction.
+Notably, Trotterization has already been implemented experimentally in superconducting quantum circuits [79], IBM Quantum’s hardware [74] and even quantum many-body platform [77] to generate dissipative dynamics, such as the dissipative TFIM.
+Moreover, tensor network methods offer promising avenues for extending Trotterization to more complex open quantum systems [72].
+We first consider the case where the original dynamics is absent [$\mathcal{L}(\rho)=0$] to gain insight.
+In this case, the generator corresponding to the reset channel simply reads
+$\dot{\rho}(t)=\mathcal{R}[\rho(t)]=r[\rho_{\delta}-\rho(t)].$ (24)
+Thus, a direct integration shows that applying the reset channel for a time $t_{s}$ maps an initial state $\rho_{0}$ to a final state $\rho(t_{s})$ according to
+$\rho(t_{s})=e^{\mathcal{R}t_{s}}[\rho_{0}]=p\rho_{0}+(1-p)\rho_{\delta},\quad p:=e^{-rt_{s}}.$ (25)
+That is, a reset channel with rate $r=-\ln p/t_{s}$ is effectively equivalent to a state preparation protocol for $\rho_{\delta}$, with success probability $1-p$.
+$p$ is a classical probability generated beforehand.
+One simply flips a biased coin (or, equivalently, performs any local measurement that yields the desired classical randomness): with probability $1-p$ the system is driven into $\rho_{\delta}$ by any suitable operations and with the probability $p$ it is left untouched.
+Since there are no restrictions on the preparation method, once $\rho_{\delta}$ is locally preparable (e.g.
+a separable state), this operation can be implemented purely with local operations, even for many-body systems.
+Specifically, when $\rho_{\delta}=\mathbb{I}/d$, Eq.
+(25) is a depolarizing channel, which can be directly realized by the 2-design [80] without flipping a coin.
+When $\mathcal{L}\neq 0$, we apply the Lie-Trotter formula [81]
+$e^{(\mathcal{L}+\mathcal{R})t}=\lim_{n\to\infty}(e^{\mathcal{R}\frac{t}{n}}e^{\mathcal{L}\frac{t}{n}})^{n}.$ (26)
+to realize the protocol. With this formula, the reset protocol during $[0,t_{s}]$ can be approximated as
+$e^{(\mathcal{L}+\mathcal{R})t_{s}}[\rho_{0}]\approx(e^{\mathcal{R}\delta t}e^{\mathcal{L}\delta t})^{n}[\rho_{0}],$ (27)
+where $\delta t:=t_{s}/n\ll 1$. Using Eq. (25), we have
+$e^{\mathcal{R}\delta t}[\rho]\approx(1-p_{s})\rho+p_{s}\rho_{\delta},$ (28)
+which can be interpreted as a state preparation mapping from $\rho$ to $\rho_{\delta}$, with success probability $p_{s}:=r\delta t$.
+Practically, $r$ is determined by $p_{s}$ and $\delta t$, whose values are set by experimentalists.
+Thus, the reset protocol may be achieved experimentally by performing the aforementioned state preparation operations stroboscopically: at each discrete time point $t_{i}=i\delta t,\ (i=1,...,n)$ within $[0,t_{s}]$, one applies the state preparation mapping from $\rho(t_{i})$ to $\rho_{\delta}$ with a given probability $p_{s}$.
+There are no constraints on the preparation time $\tau_{\rm prep}$ for $\rho_{\delta}$, but on average it adds a cost of $np_{s}\tau_{\rm prep}$ to the tailored relaxation timescale.
+Therefore, it is desirable to choose a $\rho_{\delta}$ that can be prepared efficiently, e.g.
+the maximally mixed state in our TFIM example.
+This state is convenient to prepare: coupling to an infinite-temperature bath can achieve it in $\mathcal{O}(\ln N)$ time [82] with minimal resources in a $N$-body system—typically negligible compared to the system’s intrinsic relaxation time.
+Local control methods could in principle realize it in constant time, though with resources scaling with $N$.
+This Trotterization-based method admits a natural interpretation at the level of stochastic trajectories: for a Poissonian reset process with rate $r$, the system is stochastically reset to the target state with probability $r\delta t$, in each small interval $\delta t$.
+The approximation error is of order $\mathcal{O}(t_{s}^{2}/n)$ [83, 84, 73]. Explicitly, we establish a rigorous upper bound [59]:
+$||e^{(\mathcal{L}+\mathcal{R})t_{s}}-(e^{\mathcal{R}\delta t}e^{\mathcal{L}\delta t})^{n}||\leq\frac{t_{s}^{2}}{2n}||[\mathcal{L,R}]||,$ (29)
+where |||||$\cdot$||| | ⋅ | | denotes any norm that is contractive under Lindbladian evolution, such as the trace norm or the diamond norm.
+This bound holds for arbitrary Lindbladian $\mathcal{L}$ and $\mathcal{R}$.
+In our specific case,
+$[\mathcal{L,R}]\rho:=[\mathcal{LR}-\mathcal{RL}]\rho=r[\mathcal{L}(\rho_{\delta})-\rho_{\delta}].$ (30)
+Hence, a large $n$ is not required when $t_{s}\ll 1$, $rt_{s}\ll 1$ or $\rho_{\delta}$ is close to $\rho_{\rm ss}$.
+A small $t_{s}$ with moderate $r$ offers a practical regime that may yield substantial acceleration and simultaneously reduce the required number of Trotter steps.
+Additionally, higher-order approximations could be used to reduce the error for fixed $n$.
+For instance, the second-order Suzuki-Trotter formula $e^{(\mathcal{L}+\mathcal{R})t_{s}}\approx(e^{\mathcal{L}\delta t/2}e^{\mathcal{R}\delta t}e^{\mathcal{L}\delta t/2})^{n}$ suppresses the error to $\mathcal{O}(t_{s}^{3}/n^{2})$ [83, 73].
+Dephasing noise can accelerate the relaxation of the TFIM.—Here, we add dephasing noise along a single axis (typically the $z$-axis) to the TFIM.
+The corresponding jump operators are $L^{(\phi)}_{i}=\sqrt{\gamma_{\phi}}\sigma_{i}^{z},\quad i=1,\dots,N$.
+This contributes to the total Lindbladian via an extra term
+$\mathcal{L}_{\phi}[\rho]=\gamma_{\phi}\sum_{i=1}^{N}\left(\sigma_{i}^{z}\rho\sigma_{i}^{z}-\rho\right),$ (31)
+which could be interpreted as a partial and local reset channel (only the $z$-axis is affected, and only local one-body jump operators are involoved).
+As shown in Fig.
+3, the relaxation is accelerated significantly.
+Figure 3: Dephasing noise induce acceleration of relaxation in a 5-site TFIM.
+Parameters: (a) $J=1.0,\ g=2.0,\ \gamma=0.5,\ \beta=0.1J,\ t_{s}=0.2\tau_{2}$ (b) $J=1.0,\ g=2.0,\ \gamma=0.5,\ \beta=0.1J,\ t_{s}=0.8\tau_{2}$.
+A theoretical argument supporting the finding is provided in [59].

### v4

@@
 Slow relaxation processes spanning widely separated timescales pose fundamental challenges for probing steady-state properties and engineering functional quantum systems, such as quantum heat engines and quantum computing devices.
-We introduce a protocol that enables exponential acceleration of relaxation in general Markovian open quantum systems by temporarily coupling the system to a reset channel, inspired by the quantum Mpemba effect.
+We introduce a protocol that enables significant acceleration of relaxation in general Markovian open quantum systems by temporarily coupling the system to a reset channel, inspired by the Mpemba effect.
 Crucially, this acceleration persists even when the slowest decaying Lindbladian modes form complex-conjugate pairs.
-Unlike previous approaches, which typically target a single mode, our protocol eliminates multiple relaxation modes simultaneously.
-Specifically, using a simple depolarizing channel, we show that the relaxation of a dissipative transverse-field Ising chain can be significantly accelerated, underscoring the experimental feasibility of the protocol.
-This framework provides a versatile tool for controlling quantum relaxation timescales, with broad implications for quantum thermodynamics, computation, and state preparation.
+Unlike previous approaches, which typically target a single mode, our protocol may suppress multiple relaxation modes simultaneously.
+This framework provides a versatile and experimentally feasible tool for controlling relaxation timescales, with broad implications for quantum thermodynamics, computation, and state preparation.
 Introduction.—Relaxation processes in open quantum systems, wherein a driven or interacting system settles into an equilibrium or a nonequilibrium steady state, are of fundamental importance in nonequilibrium physics and for practical quantum technologies.
 The timescales of such relaxations often determine the feasibility and performance of quantum devices.
-For example, in quantum thermodynamics, the power output of heat engines or transport junctions is limited by how rapidly the working substance reaches its steady state in each cycle [1, 2, 3].
+For example, in quantum thermodynamics, the power output of cyclic heat engines or transport junctions can be limited by how rapidly the working substance relaxes to its steady state in each cycle [1, 2].
 In addition, fast relaxation facilitates ground state laser cooling [3, 4, 5] and the reliable preparation of quantum states [6].
@@
 However, existing quantum implementations typically impose strict constraints, such as assuming pure initial states or requiring the second-largest eigenvalue of the Lindbladian to be real.
-Moreover, careful initial-state design is a delicate task, since it relies on detailed knowledge of the original initial state and the system dynamics, which is typically not known a prior in complex quantum systems.
+Moreover, careful initial-state design is delicate, since it relies on fine-tuning of control parameters and detailed knowledge of the initial state and system dynamics, which are typically not known a priori in complex quantum systems.
 These issues limit their applicability in general settings.
 Thus, a general and practical method for inducing faster relaxation in open quantum systems, which is robust with respect to different initial states, remains elusive.
-In this study, we address these challenges by proposing an experimentally feasible protocol that can both accelerate and decelerate general quantum relaxation processes from any initial state, by resetting the system to a specified (possibly mixed) target state [45, 46, 47].
+In this study, we contribute to addressing these challenges by proposing an experimentally feasible protocol that can both accelerate and decelerate general quantum relaxation processes from any initial state, by resetting the system to a specified (possibly mixed) target state [44, 45, 46].
 Here, this type of quantum reset operation can be realized through engineered dissipation, i.e., quantum reservoir engineering [47, 48, 8, 49], and should be distinguished from the conventional reset typically used for qubit initialization in quantum computation [50, 51, 52, 53, 54, 55, 56, 57].
 In our protocol, we introduce a finite-duration reset phase into the open-system dynamics by temporarily coupling the system to a reset channel.
-We show that this protocol can exponentially accelerate relaxation, even when starting from mixed initial states and in the presence of complex decay modes.
-Remarkably, our method allows for the simultaneous elimination of multiple relaxation modes—a capability absent in previous approaches.
+We show that this protocol can significantly accelerate relaxation, even when starting from mixed initial states and in the presence of complex decay modes.
+Remarkably, our method allows for the simultaneous suppression of multiple relaxation modes—a capability absent in previous approaches.
 Moreover, selective application of the protocol enables the system of interest to remain in metastable states for extended periods, offering a systematic route to suppress relaxation.
-Setup.—We consider a general open quantum system defined on a Hilbert space $\mathcal{H}$ of dimension $d$, whose dynamics in the interaction picture are governed by the Lindblad-Gorini-Kossakowski-Sudarshan (LGKS) master equation $d\rho/dt=\mathcal{L}(\rho),$ where the Lindbladian superoperator $\mathcal{L}$ is given by
+Setup.—We consider a general Markovian open quantum system defined on a Hilbert space $\mathcal{H}$ of dimension $d$, whose dynamics are governed by the Lindblad-Gorini-Kossakowski-Sudarshan (LGKS) master equation $d\rho/dt=\mathcal{L}(\rho),$ where the Lindbladian superoperator $\mathcal{L}$ is given by
 $\mathcal{L}(\rho)\equiv-i[H,\rho]+\sum_{i}\left[J_{i}\rho J_{i}^{\dagger}-\frac{1}{2}\{J_{i}^{\dagger}J_{i},\rho\}\right].$ (1)
@@
 The evolution can be analyzed via the spectral decomposition of $\mathcal{L}$, whose eigenvalues are denoted $\lambda_{k}$ and ordered such that $0\geq\text{Re}(\lambda_{k})\geq\text{Re}(\lambda_{k+1})$.
-Assuming that $\mathcal{L}$ is diagonalizable, let $R_{k}$ and $L_{k}$ denote the corresponding right and left eigenmatrices, satisfying $\mathcal{L}(R_{k})=\lambda_{k}R_{k}$ and $\mathcal{L^{\dagger}}(L_{k})=\lambda_{k}^{\star}L_{k},\ k=1,...,d^{2}$, with biorthogonal normalization
+Assuming that $\mathcal{L}$ is diagonalizable, let $R_{k}$ and $L_{k}$ denote the corresponding right and left eigenmatrices, satisfying $\mathcal{L}(R_{k})=\lambda_{k}R_{k}$ and $\mathcal{L^{\dagger}}(L_{k})=\lambda_{k}^{\star}L_{k},\ k=1,...,d^{2}$, with biorthogonal normalization [58]
 $\text{Tr}(L_{k}^{\dagger}R_{h})=\delta_{kh}.$ (2)
@@
 Given an initial state $\rho_{0}$, the system state at time $t$ evolves as
-$\rho(t)=e^{t\mathcal{L}}[\rho_{0}]=\rho_{\rm ss}+\sum_{k=2}^{d^{2}}e^{\lambda_{k}t}c_{k}R_{k},$ (3)
+$\rho(t)=e^{t\mathcal{L}}[\rho_{0}]=\rho_{\rm ss}+\sum_{k=2}^{d^{2}}c_{k}e^{\lambda_{k}t}R_{k},$ (3)
 where $c_{k}\equiv\text{Tr}(L_{k}^{\dagger}\rho_{0})$.
 The unique stationary state of the open quantum system $\rho_{\rm ss}$ is given by $\rho_{\rm ss}=\lim_{t\rightarrow\infty}\rho(t)=R_{1}$, assuming $\lambda_{1}=0$ is non-degenerate.
-In the long-time limit, the relaxation is dominated by the slowest decaying mode, and the deviation from stationarity obeys $||\rho(t)-\rho_{\rm ss}||\sim\exp\left(-|\text{Re}\lambda_{2}|t\right),$ so the characteristic relaxation timescale is given by $\tau_{{\rm{\rm rel}}}\sim\frac{1}{|\text{Re}(\lambda_{2})|}$.
-For clarity of exposition, we first consider the simple case where the second eigenvalue $\lambda_{2}$ is real and non-degenerate, as in [10].
-Later, we will generalize the discussion to include cases where the subleading eigenvalues form complex conjugate pairs.
-Reset protocol to accelerate relaxation processes.—Recognizing that the slowest decaying mode governs the relaxation timescale, we propose a quantum reset protocol that suppresses its excitation from a general initial state via a finite-duration reset phase, thereby enabling exponential acceleration of relaxation.
+In the long-time limit, the relaxation is dominated by the slowest decaying mode, and the deviation from stationarity obeys $\rho(t)-\rho_{\rm ss}||\sim\exp\left(-|\text{Re}\lambda_{2}|t\right)$.
+Reset protocol to accelerate or decelerate relaxation processes.—Recognizing that the slowest decaying mode governs the relaxation timescale, we propose a quantum reset protocol that can suppress or promote its excitation from a general initial state via a finite-duration reset phase, thereby enabling exponential acceleration of relaxation.
 Specifically, we let the system evolve under modified dynamics for a finite time $t_{s}$, during which it is stochastically reset to a chosen state $\rho_{\delta}$ at random times, with events following a Poisson process of rate $r$.
@@
 That is, the system density matrix $\rho(t)$ evolves under the reset channel for $t\in[0,t_{s}]$ and decouples from the channel for $t>t_{s}$.
-The modified dynamics during the reset phase is governed by a new Lindbladian $\mathcal{L}_{r}$ [45], defined as
-$\mathcal{L}_{r}(\rho)=\mathcal{L}(\rho)+r\text{Tr}(\rho)\rho_{\delta}-r\rho.$ (4)
-Expressing $\rho_{\delta}=\sum_{\alpha}p_{\alpha}|\psi_{\alpha}\rangle\langle\psi_{\alpha}|$, $\mathcal{L}_{r}$ is equivalent to introducing jump operators $J_{i}^{r,\alpha}=\sqrt{rp_{\alpha}}|\psi_{\alpha}\rangle\langle\phi_{i}|$ into $\mathcal{L}$, where $|\phi_{i}\rangle$ form an auxiliary complete orthonormal basis (see End Matter).
+The modified dynamics during the reset phase $t\in[0,t_{s}]$ is governed by a new Lindbladian $\mathcal{L}_{r}$ [44], defined as
+$\mathcal{L}_{r}(\rho):=\mathcal{L}(\rho)+\mathcal{D}_{r}(\rho)=\mathcal{L}(\rho)+r\text{Tr}(\rho)\rho_{\delta}-r\rho,$ (4)
+where $\mathcal{D}_{r}(\rho)=r[\text{Tr}(\rho)\,\rho_{\delta}-\rho]$.
+Expressing $\rho_{\delta}=\sum_{\alpha}p_{\alpha}|\psi_{\alpha}\rangle\langle\psi_{\alpha}|$, $\mathcal{L}_{r}$ is equivalent to introducing jump operators $J_{i,\alpha}^{r}=\sqrt{rp_{\alpha}}|\psi_{\alpha}\rangle\langle\phi_{i}|$ into $\mathcal{L}$, where $|\phi_{i}\rangle$ form an auxiliary complete orthonormal basis (see End Matter).
 The term $r\text{Tr}(\rho)\rho_{\delta}$ contributes only to the stationary mode and vanishes on all traceless eigenmodes $R_{i}$ with $i\geq 2$, which satisfy $\mathrm{Tr}(R_{i})=0$ due to $\mathrm{Tr}(L_{1}^{\dagger}R_{i})=0$ and $L_{1}^{\dagger}=\mathbb{I}$ [cf.
@@
 $\mathcal{L}_{r}(R_{i})=(\lambda_{i}-r)R_{i},\ i\in\{2,...,d^{2}\}.$ (5)
-This induces a uniform spectral shift by $-r$ for all non-stationary eigenmodes, leaving the eigenmatrices ($R_{i}$, $i\geq 2$) themselves unchanged.
+This induces a uniform spectral shift by r-r for all non-stationary eigenmodes, leaving the eigenmatrices ($R_{i}$, $i\geq 2$) themselves unchanged.
 In contrast, the stationary mode $R_{1}$ retains eigenvalue zero but is modified under reset.
@@
 $\rho^{r}(t)$ $=e^{t\mathcal{L}_{r}}\rho_{\rm ss}+\sum_{k=2}^{d^{2}}c_{k}e^{(\lambda_{k}-r)t}R_{k}$
-$:=\rho_{\rm ss}+\sum_{k}c_{k}^{r}(t)e^{\lambda_{k}t}R_{k},$ (6)
+$:=\rho_{\rm ss}+\sum_{k}c_{k}^{r}(t)e^{\lambda_{k}t}R_{k},\ (0\leq t\leq t_{s})$ (6)
 where $c_{k}^{r}(t)$ are time-dependent mode amplitudes. Explicitly, these modified coefficients are given by (End Matter)
 $c_{k}^{r}(t)\equiv\left[c_{k}-\frac{r\cdot d_{k}}{r-\lambda_{k}}\right]e^{-rt}+\frac{r\cdot d_{k}}{r-\lambda_{k}}e^{-\lambda_{k}t},$ (7)
-where $d_{k}=\text{Tr}(L_{k}^{\dagger}\rho_{\delta})$.
-By properly choosing an optimal $t^{*}_{s}$, we can set $c_{2}^{r}(t^{*}_{s})=0$, thereby eliminating the overlap with the slowest decaying mode.
-Solving $c_{2}^{r}(t_{s})=0$ yields
-$t^{*}_{s}(r)=\frac{1}{r-\lambda_{2}}\ln\left[1-\frac{c_{2}(r-\lambda_{2})}{r\cdot d_{2}}\right].$ (8)
-By an appropriate choice of $\rho_{\delta}$ such that $c_{2}/d_{2}\leq 0$, $t^{*}_{s}(r)\geq 0$ can be assured.
-After the reset phase ($t>t_{s}$), the system returns to original Lindbladian evolution, now without the contribution from the slowest mode.
-The full evolution reads:
-$\rho^{r}(t)=$ $\rho_{\rm ss}+\sum_{k}c_{k}^{r}(t)e^{\lambda_{k}t}\Theta(t_{s}-t)R_{k}$
-$+\sum_{k}c_{k}^{r}(t_{s})e^{\lambda_{k}t}\Theta(t-t_{s})R_{k},$ (9)
-where $\Theta(x)$ is the Heavside step function. This protocol with an optimal $t_{s}=t^{*}_{s}$ reduces the relaxation time to
-$\tau_{{\rm rel}}\sim t^{*}_{s}+\frac{1}{|{\rm Re}(\lambda_{3})|},$ (10)
-which is generically shorter than the unperturbed timescale $\tau_{{\rm{\rm rel}}}\sim\frac{1}{|\text{Re}(\lambda_{2})|}$.
+where $d_{k}\equiv\text{Tr}(L_{k}^{\dagger}\rho_{\delta})$ is the overlap coefficient of $\rho_{\delta}$.
+Notably, Eq.
+(Accelerating Quantum Relaxation via Temporary Reset: A Mpemba-Inspired Approach) can be generalized to the dynamics of observables [59], which may be more convenient to measure.
+After the reset phase ($t>t_{s}$), the system returns to the original Lindbladian evolution. The dynamics for $t>t_{s}$ then reads:
+$\rho^{r}(t)=\rho_{\rm ss}+\sum_{k}c_{k}^{r}(t_{s})e^{\lambda_{k}t}R_{k}.\quad(t>t_{s})$ (8)
+Importantly, the modified spectral decomposition, Eq.
+(8), takes the same form as the original one, Eq.
+(3), which allows us to directly compare the relaxation dynamics with and without reset.
+From this comparison, it is clear that for $t>t_{s}$ our protocol is equivalent to the original Lindbladian evolution, but with different overlap coefficients,
+$\rho(t)=\rho_{\rm ss}+\sum_{k=2}^{d^{2}}c^{\prime}_{k}e^{\lambda_{k}t}R_{k}.\quad(t>t_{s})$ (9)
+Here, $c_{k}^{\prime}\equiv c^{r}_{k}(t_{s})$ are determined by the reset protocol.
+Thus, our protocol controls relaxation dynamics by tuning the overlap coefficients $c_{k}$, without requiring any special preparation of the initial state.
+By choosing an appropriate $\rho_{\delta}$ and varying $r$ and $t_{s}$, one can tune the amplitudes of multiple relaxation modes.
+In particular, ensuring $|c_{2}^{r}(t_{s})|<|c_{2}|$ ($|c_{2}|$) suppresses (enhances) the dominant mode, which is sufficient for acceleration (deceleration), ultimately bringing the system closer to (farther from) stationarity [19, 32].
+Specifically, the dominant mode is fully eliminated and exponential speed-up occurs when $|c_{2}^{r}(t_{s})|=0$; this, however, requires fine-tuning and prior knowledge and is therefore challenging to realize in large systems.
+We thus focus on the weaker but practically relevant scenario $|c_{2}^{r}(t_{s})|<|c_{2}|$.
+Sufficient conditions for suppression or promotion of relaxation modes.—We derive a sufficient condition [59],
+${\rm Re}(c_{2}^{*}d_{2})<|c_{2}|^{2},$ (10)
+under which there always exists a threshold $t_{c}>0$ such that, for any $r>0$ and any $t_{s}\in(0,t_{c}]$, one has $|c_{2}^{r}(t_{s})|<|c_{2}|$.
+That is, the relaxation is accelerated for all $t_{s}$ in this range.
+If the condition is violated, there exists another threshold $t_{c}^{\prime}>0$ with $|c_{2}^{r}(t_{s})|\geq|c_{2}|$ for $t_{s}\in(0,t_{c}^{\prime}]$, leading to deceleration.
+However, Eq.
+(10) is not necessary for acceleration, since one may still have $|c_{2}^{r}(t_{s})|<|c_{2}|$ for $t_{s}>t_{c}^{\prime}$, when it is not satisfied.
+Several remarks are in order. First, the condition applies both when $\lambda_{2}$ is real and when it forms a complex-conjugate pair. In the latter case, the long-time dynamics takes the form
+$\rho(t)=\rho_{\rm ss}+e^{\lambda_{2}t}c_{2}R_{2}+e^{\lambda_{2}^{*}t}c_{2}^{*}R_{2}^{\dagger}+\mathcal{O}(e^{\lambda_{3}t}),$ (11)
+where $c_{2}=\text{Tr}(L_{2}^{\dagger}\rho_{0})$ and $c_{2}^{*}=\text{Tr}(L_{2}\rho_{0})$ is its complex conjugate.
+Most existing strategies cannot simultaneously suppress both components of such complex relaxation modes.
+Our protocol, however, naturally addresses this limitation.
+The modified coefficients under reset are
+$c_{2}^{r}(t)$ $=\left[c_{2}-\frac{r\,d_{2}}{r-\lambda_{2}}\right]e^{-rt}+\frac{r\,d_{2}}{r-\lambda_{2}}e^{-\lambda_{2}t},$ (12a)
+$c_{2}^{r,*}(t)$ $=\left[c_{2}^{*}-\frac{r\,d_{2}^{*}}{r-\lambda_{2}^{*}}\right]e^{-rt}+\frac{r\,d_{2}^{*}}{r-\lambda_{2}^{*}}e^{-\lambda_{2}^{*}t}.$ (12b)
+These coefficients remain complex conjugates for all $t$, i.e., $c_{2}^{r}(t)]^{*}=c_{2}^{r,*}(t)$.
 Eq.
-(9) can be generalized to the dynamics of observables [59], which may be more convenient to measure.
-Our protocol also enables controlled deceleration of relaxation.
-By choosing $\rho_{\delta}$ and $t_{s}$ to amplify the magnitude of $c_{k}^{r}(t)$, one can prolong the time during which the system remains away from the steady state.
-This capability could be useful in quantum metrology and computation, where long-lived metastable states play a functional role.
-Our reset protocol thus provides a general framework for controlling relaxation processes in open quantum systems.
-Extension to complex-conjugate relaxation modes.—We now consider the more general case where the dominant relaxation modes form a complex-conjugate pair.
-In this case, the long-time dynamics of the density matrix takes the form
-$\rho(t)=\rho_{\rm ss}+e^{\lambda_{2}t}c_{2}R_{2}+e^{\lambda_{2}^{\dagger}t}c_{2}^{*}R_{2}^{\dagger}+\mathcal{O}(e^{\lambda_{3}t}),$ (11)
-where $c_{2}=\text{Tr}(L_{2}^{\dagger}\rho_{0})$ and $c_{2}^{*}=\text{Tr}(L_{2}\rho_{0})$ is its complex conjugate.
-Most known strategies are unable to simultaneously suppress both components of such complex relaxation modes.
-Our reset protocol, however, naturally addresses this limitation. The modified coefficients under quantum reset are given by
-$c_{2}^{r}(t)=\left[c_{2}-\frac{r\cdot d_{2}}{r-\lambda_{2}}\right]e^{-rt}+\frac{r\cdot d_{2}}{r-\lambda_{2}}e^{-\lambda_{2}t}$ , (12a)
-$c_{2}^{r,*}(t)=\left[c_{2}^{*}-\frac{r\cdot d_{2}^{*}}{r-\lambda_{2}^{*}}\right]e^{-rt}+\frac{r\cdot d_{2}^{*}}{r-\lambda_{2}^{*}}e^{-\lambda_{2}^{*}t}$ (12b)
-Importantly, we observe that these coefficients remain complex conjugates of each other for all $t$, i.e., $[c_{2}^{r}(t)]^{*}=c_{2}^{r,*}(t)$.
-Therefore, eliminating the entire complex mode reduces to a single complex condition:
-$c_{2}^{r}(t_{s})=0\Leftrightarrow\text{Re}[c_{2}^{r}(t_{s})]=\text{Im}[c_{2}^{r}(t_{s})]=0,$ (13)
-because $c_{2}^{r}(t_{s})=0\Rightarrow c_{2}^{r,*}(t_{s})=[c_{2}^{r}(t_{s})]^{*}=0$.
-In some cases, this condition can be exactly satisfied by tuning $r$ and $t_{s}$.
-Although Eq.
-(13) does not always admit an exact solution, in practice, it suffices to reduce $|c_{2}^{r}(t_{s})|$ to an exponentially small value ($\text{Re}[c_{2}^{r}(t_{s})]\ll 1,\ \text{Im}[c_{2}^{r}(t_{s})]\ll 1$), which still leads to exponential acceleration of relaxation.
-Additionally, we show in [59] that, if the condition
-${\rm Re}(c_{2}^{*}d_{2})\leq|c_{2}|^{2}$ (14)
-is satisfied, $|c_{2}^{r}(t_{s})|<|c_{2}|$ can always be achieved for any positive $t_{s}$ that below a certain threshold.
-This is a weak condition that be readily satisfied for many choices of $\rho_{\delta}$.
-It is strictly weaker than the commonly satisfied condition $c_{2}/d_{2}\leq 0$, which, in the case where $\lambda_{2}$ is real, guarantees the existence of a $t_{s}^{*}$ such that $c^{r}_{2}(t_{s}^{*})=0$.
-Eq.
-(14) thus provides a sufficient condition for reset-induced Mpemba-type acceleration [20].
-It implies that fine-tuning of either $\rho_{\delta}$ or $t_{s}$ is unnecessary for achieving acceleration.
-We illustrate this behavior concretely in the following examples.
+(10) then ensures that $|c_{2}^{r,*}(t_{s})|=|c_{2}^{r}(t_{s})|\leq|c_{2}|=|c_{2}^{*}|$ for a moderate $t_{s}$.
+Second, Eq.
+(10) is a weak condition satisfied by many choices of $\rho_{\delta}$.
+In particular, it does not require that $\rho_{\delta}$ be closer to the steady state than $\rho_{0}$, nor that the overlap coefficient $|d_{2}|$ be smaller than $|c_{2}|$.
+For example, when $\lambda_{2}$ is real, if ${\rm Re}(d_{2})$ and $c_{2}$ have opposite signs, the condition is satisfied even when $|d_{2}|$ is arbitrarily large.
+In Example 2 we show that the condition is typically satisfied for a large fraction of randomly chosen initial states.
+Third, neither $r$ nor $t_{s}$ needs to be precisely tuned in advance to achieve acceleration or deceleration.
+Provided that Eq.
+(10) holds, any positive $r$ and moderately small $t_{s}$ are sufficient for acceleration, and $t_{s}$ can be further optimized a posteriori based on the observed dynamics.
+Our protocol therefore provides a general and robust framework for controlling relaxation in open quantum systems, without the need for fine-tuning.
+Finally, it is straightforward to see that
+${\rm Re}(c_{k}^{*}d_{k})<|c_{k}|^{2}$ (13)
+is a sufficient condition for suppression of the $k$-th mode.
+Since multiple such conditions for different $k$ can be satisfied simultaneously, our protocol allows simultaneous suppression (or promotion) of multiple relaxation modes.
+In [59] we provide numerical evidence that multiple conditions can indeed be satisfied in parallel, using the model and protocol of Example 2.
+So far we have focused on practical scenarios where fine-tuning is not possible and neither the system dynamics nor the initial state are known a priori.
+If these constraints are lifted, as in most previous studies, our protocol can fully eliminate the dominant relaxation mode ($|c_{2}^{r}(t_{s})|=0$) by choosing an optimal $t_{s}=t_{s}^{*}$.
+When $\lambda_{2}$ is real, the exact expression is [59, 24]
+$t^{*}_{s}=\frac{1}{r-\lambda_{2}}\ln\!\left(1-\frac{c_{2}(r-\lambda_{2})}{rd_{2}}\right).$ (14)
+A sufficient condition ensuring that $t^{*}_{s}\geq 0$ is $c_{2}/d_{2}\leq 0$. Further details, including the case of complex-conjugate pairs, are given in [59].
 Example 1.—To illustrate our protocol, we first consider a minimal example: a two-level quantum system, which may represent a qubit or a spin-$1/2$ particle.
@@
 The system is coupled to a thermal environment through the Lindblad jump operators $J_{0}=\sqrt{\gamma_{0}}\sigma^{+}$ and $J_{1}=\sqrt{\gamma_{1}}\sigma_{-}$.
-Here, $\gamma_{0}$ ($\gamma_{1}$) is the transition rate from state $|0\rangle$ to state $|1\rangle$ (or converse).
+Here, $\gamma_{0}$ ($\gamma_{1}$) is the transition rate from state $|0\rangle$ to state $|1\rangle$ (or vice versa).
 Assuming the rates satisfy detailed balance with inverse temperature $\gamma_{0}e^{-\beta_{\text{env}}E_{0}}=\gamma_{1}e^{-\beta_{\text{env}}E_{1}}$, i.e., $\gamma_{0}=\gamma_{1}e^{-\beta_{\text{env}}E},$ the system evolves to a unique equilibrium state $\rho_{\rm eq}=\frac{e^{-\beta_{\text{env}}H}}{\text{Tr}(e^{-\beta_{\text{env}}H})}$ irrespective of the initial condition.
@@
 where the off-diagonal terms $ke^{\pm i\phi}$ quantify initial quantum coherence.
-In the absence of intrinsic transitions ($\Omega=0$), the Hamiltonian is diagonal and the relaxation process reduces to pure decoherence.
-In this case, all eigenmodes and eigenvalues of the Lindbladian can be computed analytically [59].
-When we apply our reset protocol with the reset state $\rho_{\delta}=|0\rangle\langle 0|$, corresponding to the ground state to ensure $t_{s}(r)\geq 0$, the two slowest decaying modes form a complex-conjugate pair:
+When $\Omega=0$, all eigenmodes and eigenvalues of the Lindbladian can be computed analytically [59].
+When we apply our reset protocol with the reset state $\rho_{\delta}=|0\rangle\langle 0|$, corresponding to the ground state, the two slowest decaying modes form a complex-conjugate pair:
 $c_{2}^{r}(t)=e^{-rt+i\phi}k,\quad c_{2}^{r,*}(t)=e^{-rt-i\phi}k.$ (15)
-This overlap cannot be entirely eliminated in finite time, but it decays exponentially fast, leading to a relaxation timescale
-$\tau_{{\rm rel}}\sim\frac{1}{r}+\frac{1}{|\text{Re}(\lambda_{3})|},$ (16)
-where $1/r$ characterizes the reset protocol timescale, and $1/|\text{Re}(\lambda_{3})|$ represents the accelerated relaxation timescale with the slowest mode eliminated.
-Clearly, $\tau_{{\rm rel}}<1/|{\rm Re}(\lambda_{2})|$ for sufficiently large $r$.
-By appropriately tuning the reset rate, the relaxation timescale can approach $1/|\rm Re(\lambda_{3})|$ arbitrarily closely.
-Using Eq.
-(9), the full dynamics under our reset protocol can thus be determined.
-Since the condition $c_{2}^{r}(t_{s})=0$ cannot be satisfied exactly in this case—though $|c_{2}^{r}(t_{s})|$ can be made exponentially small by choosing $t_{s}\sim 1/r$—we instead define $t_{s}$ such that a faster decaying mode is suppressed.
-Specifically, solving $c_{4}^{r}(t_{s})=0$ gives
-$t_{s}=\frac{\ln\left[1+\frac{\gamma_{1}k_{\text{tot}}}{r}\left(\frac{1}{\gamma_{0}+\gamma_{1}e^{(\beta_{0}-\beta_{\text{env}})E}}-\frac{1}{\gamma_{0}+\gamma_{1}}\right)\right]}{k_{\text{tot}}},$ (17)
-where the total transition rate is $k_{\text{tot}}=r+\gamma_{0}+\gamma_{1}$.
-This choice fully eliminates the fastest decaying relaxation mode.
-In addition, for a sufficiently large reset rate $r$, the slowest mode’s amplitude becomes exponentially small after the reset phase.
+These overlaps are exponentially suppressed compared with the original overlap, i.e.
+$c_{2}^{r}(t)=e^{-rt}c_{2}$ and $c_{2}^{r,*}(t)=e^{-rt}c_{2}^{*}$.
+Clearly, the relaxation is accelerated for any $r$ and $t_{s}$.
+Since any $t_{s}$ can ensure acceleration, we choose $t_{s}$ such that a faster decaying mode is eliminated.
+Specifically, solving $c_{4}^{r}(t_{s})=0$ gives such a $t_{s}$, whose expression is provided in [59].
+For a sufficiently large $r$, the slowest mode’s amplitude becomes exponentially small after the reset phase.
 Alternative selections for $t_{s}$ can also achieve accelerated relaxation, indicating that detailed dynamics knowledge is not essential.
@@
 We characterize the distance between the transient state $\rho(t)$ and the stationary state $\rho_{\rm ss}$ using the standard trace distance $D[\rho(t)|\rho_{\rm ss}]:=\text{Tr}|\rho(t)-\rho_{\rm ss}|/2$ where $|A|:=\sqrt{A^{\dagger}A}$.
-Other measures, such as the $L_{\infty}$ norm $D_{\infty}[\rho(t)|\rho_{\rm ss}]:=\max_{i}|\lambda_{i}|$ [$\lambda_{i}$ is the $i$-th eigenvalue of $\rho(t)-\rho_{\rm ss}$], yield qualitatively similar outcomes.
-We first plot $D[\rho(t)|\rho_{\rm eq}]$ over $t$ for various rate $r$ in Fig.
+Other measures, such as the $L_{\infty}$ norm $D_{\infty}[\rho(t)|\rho_{\rm ss}]:=\max_{i}|\lambda_{i}|$ $\lambda_{i}$ is the $i$-th eigenvalue of $\rho(t)-\rho_{\rm ss}$], yield qualitatively similar outcomes.
+We first plot $D[\rho(t)|\rho_{\rm eq}]$ over $t$ for various rates $r$ in Fig.
 1 (a).
-The parameters for $r>0$ cases are: initial conditions $\beta_{0}=2.0$, $k=0.32$ and $\phi=1$; The $r=0$ scenario has $\beta_{0}=3.0,$ $k=0.21$ and $\phi=1$.
-The environment is at lower temperature $\beta_{\text{env}}=4.0$.
-Although the $r>0$ cases are initially farther from equilibrium compared to the $r=0$ case, the reset protocol substantially accelerates relaxation so that the former cases reach equilibrium even faster.
+The parameters for $r>0$ cases are chosen so that they are initially farther from equilibrium compared to the $r=0$ case (see the caption of Fig.
+1).
+The reset protocol substantially accelerates relaxation so that the former cases reach stationarity faster even if they are farther from stationarity initially.
 We next plot $t_{s}$ as a function of $r$ in Fig.
@@
 1 (c) confirms accelerated relaxation for $\Omega=2$, here measured by the $L_{\infty}$ norm to reduce fluctuations in the curves.
+Notably, the steady state is not equilibrium when $\Omega>0$.
 Figure 1: Relaxation dynamics of the two-state system.
-(a) and (c): Distances between $\rho(t)$ and $\rho_{\text{eq}}$ as the function of time $t$ with different resetting rate $r$.
+(a) and (c): Distances between $\rho(t)$ and $\rho_{\text{eq}}$ as a function of time $t$ with different resetting rates $r$.
 (b) and (d): The $t_{s}$ as a function of the resetting rate $r$.
 The parameters are chosen as $\beta_{0}=2.0$, $k=0.32$ and $\phi=1$ with reset protocol and $\beta_{0}=3.0,$ $k=0.21$ and $\phi=1$ without reset.
+The environment is at lower temperature $\beta_{\text{env}}=4.0$.
+$\gamma_{1}=1.0$.
 For (a) and (b) $\Omega=0$, for (c) and (d) $\Omega=2.0$.
@@
 Here, $\rho_{\rm ss}$ is nontrivial, with nonzero entanglement.
+Moreover, as shown in [59], for a large fraction of randomly chosen pure initial states our protocol with $\rho_{\delta}=\mathbb{I}/d$ still yields substantial acceleration, further demonstrating its robustness.
 Parameters: $t_{s}=0.5\,\tau_{2}$ and $t_{s}=0.08\,\tau_{2}$ with $\tau_{2}=1/|\operatorname{Re}\lambda_{2}|$, the relaxation timescale, total duration $T=6\,\tau_{2}$, $\alpha=0.55/0.05$.
-Notably, $t_{s}$ is chosen arbitrarily, rather than being determined a prior.
+Notably, $t_{s}$ is chosen arbitrarily, rather than being determined a priori.
 The accelerated relaxation processes under different reset rate are shown in Fig.
@@
 The existence of such an effect broadens the range of usable reset states.
+We further verify in [59] that the same protocol significantly accelerates relaxation in another many-body setting, the Dicke model, again using $\rho_{\delta}=\mathbb{I}/d$.
 Some remarks are in order regarding experimental implications.
-In general, the optimal duration $t_{s}$ of the reset phase cannot be determined a priori from theoretical considerations, as the relevant dynamical details are typically unknown in complex quantum systems.
-The TFIM example shows how this challenge can be overcome: fine-tuning of $t_{s}$ is not required—any finite value shorter than the native relaxation time leads to substantial speed-up, and $t_{s}$ can be further adjusted a posteriori based on observed dynamics.
-This robustness, without requiring prior knowledge of the system dynamics and initial states, distinguishes our protocol from previous approaches.
+In practice, the precise optimal duration $t_{s}$ is generally unknown in advance, but as emphasized before, fine-tuning is unnecessary.
+The TFIM example directly supports this conclusion: any choice of $t_{s}$ shorter than the intrinsic relaxation time already leads to a clear acceleration.
 Additionally, the reset channel in the example corresponds to a depolarizing channel, which is conceptually straightforward to implement, e.g., by applying an isotropic white-noise field or coupling an infinite-temperature reservoir.
 Figure 2: Acceleration of the relaxation of TFIMs.
-Reset rate $r$ are chosen as $0,1.0,5.0,10.0,20.0$.
+Reset rates $r$ are chosen as $0,1.0,5.0,10.0,20.0$.
 (a) $N=5$ ($d=32$), $\alpha=0.55$, $t_{s}=0.08\tau_{2}$.
@@
 (d) $N=6$, $\alpha=0.05$, $t_{s}=0.50\tau_{2}$.
-The vertical gray lines indicate $t_{s}$ when the reset channel is removed.
+The vertical gray lines mark $t_{s}$ at which the reset channel is removed.
 Inset: Schematic representation of the dissipative TFIM, with each spin coupled to a thermal bath.
@@
 This suggests that our protocol may be testable and applicable on existing platforms.
+To connect our theoretical protocol more directly with experiments, we provide in [59] proposals that relate $r$ to experimentally tunable parameters, offering realistic paths towards implementation.
 Concluding remarks.—We have introduced a general framework for accelerating relaxation in open quantum systems via temporary reset, applicable to arbitrary initial states.
@@
 Overall, our results establish temporary reset as a powerful and experimentally feasible tool for controlling relaxation timescales in open quantum dynamics.
+We are grateful to anonymous referees for their valuable comments, which greatly enhance the quality of this work.
 This work is partly supported by the Innovation Program for Quantum Science and Technology (2021ZD0303306), MOST(2022YFA1303100).
@@
 The associated Lindblad term is
-$\mathcal{D}_{r}(\rho)=\sum_{i=1}^{d}\!\bigl{(}J_{i}^{r}\rho\,J_{i}^{r\dagger}-\tfrac{1}{2}\{J_{i}^{r\dagger}J_{i}^{r},\rho\}\bigr{)}.$ (19)
+$\mathcal{D}_{r}(\rho)=\sum_{i,\alpha}\!\bigl(J_{i,\alpha}^{r}\rho\,J_{i,\alpha}^{r\dagger}-\tfrac{1}{2}\{J_{i,\alpha}^{r\dagger}J_{i,\alpha}^{r},\rho\}\bigr).$ (17)
 Direct evaluation gives
@@
 Moreover, tensor network methods offer promising avenues for extending Trotterization to more complex open quantum systems [72].
-We first consider the case where the original dynamics is absent [$\mathcal{L}(\rho)=0$] to gain insight.
+We first consider the case where the original dynamics is absent $\mathcal{L}(\rho)=0$] to gain insight.
 In this case, the generator corresponding to the reset channel simply reads
@@
 $\rho(t_{s})=e^{\mathcal{R}t_{s}}[\rho_{0}]=p\rho_{0}+(1-p)\rho_{\delta},\quad p:=e^{-rt_{s}}.$ (23)
-That is, a reset channel with rate $r=-\ln p/t_{s}$ is effectively equivalent to a state preparation protocol for $\rho_{\delta}$, with success probability $1-p$.
+That is, a reset channel with rate $r=-\ln p/t_{s}$ is effectively equivalent to a state preparation protocol for $\rho_{\delta}$, with success probability 1p1-p.
 $p$ is a classical probability generated beforehand.
-One simply flips a biased coin (or, equivalently, performs any local measurement that yields the desired classical randomness): with probability $1-p$ the system is driven into $\rho_{\delta}$ by any suitable operations and with the probability $p$ it is left untouched.
+One simply flips a biased coin (or, equivalently, performs any local measurement that yields the desired classical randomness): with probability 1p1-p the system is driven into $\rho_{\delta}$ by any suitable operations and with the probability $p$ it is left untouched.
 Since there are no restrictions on the preparation method, once $\rho_{\delta}$ is locally preparable (e.g.
@@
 Specifically, when $\rho_{\delta}=\mathbb{I}/d$, Eq.
-(25) is a depolarizing channel, which can be directly realized by the 2-design [80] without flipping a coin.
+(23) is a depolarizing channel, which can be directly realized by a unitary 2-design [59, 80] without flipping a coin.
 When $\mathcal{L}\neq 0$, we apply the Lie-Trotter formula [81]
 $e^{(\mathcal{L}+\mathcal{R})t}=\lim_{n\to\infty}(e^{\mathcal{R}\frac{t}{n}}e^{\mathcal{L}\frac{t}{n}})^{n}.$ (24)
-to realize the protocol. With this formula, the reset protocol during $[0,t_{s}]$ can be approximated as
+to realize the protocol. With this formula, the reset protocol during $0,t_{s}]$ can be approximated as
 $e^{(\mathcal{L}+\mathcal{R})t_{s}}[\rho_{0}]\approx(e^{\mathcal{R}\delta t}e^{\mathcal{L}\delta t})^{n}[\rho_{0}],$ (25)
@@
 Practically, $r$ is determined by $p_{s}$ and $\delta t$, whose values are set by experimentalists.
-Thus, the reset protocol may be achieved experimentally by performing the aforementioned state preparation operations stroboscopically: at each discrete time point $t_{i}=i\delta t,\ (i=1,...,n)$ within $[0,t_{s}]$, one applies the state preparation mapping from $\rho(t_{i})$ to $\rho_{\delta}$ with a given probability $p_{s}$.
+Thus, the reset protocol may be achieved experimentally by performing the aforementioned state preparation operations stroboscopically: at each discrete time point $t_{i}=i\delta t,\ (i=1,...,n)$ within $0,t_{s}]$, one applies the state preparation mapping from $\rho(t_{i})$ to $\rho_{\delta}$ with a given probability $p_{s}$.
 There are no constraints on the preparation time $\tau_{\rm prep}$ for $\rho_{\delta}$, but on average it adds a cost of $np_{s}\tau_{\rm prep}$ to the tailored relaxation timescale.
@@
 The approximation error is of order $\mathcal{O}(t_{s}^{2}/n)$ [83, 84, 73]. Explicitly, we establish a rigorous upper bound [59]:
-$||e^{(\mathcal{L}+\mathcal{R})t_{s}}-(e^{\mathcal{R}\delta t}e^{\mathcal{L}\delta t})^{n}||\leq\frac{t_{s}^{2}}{2n}||[\mathcal{L,R}]||,$ (29)
-where |||||$\cdot$||| | ⋅ | | denotes any norm that is contractive under Lindbladian evolution, such as the trace norm or the diamond norm.
+$e^{(\mathcal{L}+\mathcal{R})t_{s}}-(e^{\mathcal{R}\delta t}e^{\mathcal{L}\delta t})^{n}||\leq\frac{t_{s}^{2}}{2n}$||[$\mathcal{L,R}$]$|$, (27)
+where $||\cdot||$ denotes any norm that is contractive under Lindbladian evolution, such as the trace norm or the diamond norm.
 This bound holds for arbitrary Lindbladian $\mathcal{L}$ and $\mathcal{R}$.
 In our specific case,
-$[\mathcal{L,R}]\rho:=[\mathcal{LR}-\mathcal{RL}]\rho=r[\mathcal{L}(\rho_{\delta})-\rho_{\delta}].$ (30)
+$\mathcal{L,R}]\rho:=[\mathcal{LR}-\mathcal{RL}]\rho=r{\rm Tr}(\rho)\mathcal{L}(\rho_{\delta}),$ (28)
+where ${\rm Tr}(\mathcal{L}(\rho))=0$ is used.
 Hence, a large $n$ is not required when $t_{s}\ll 1$, $rt_{s}\ll 1$ or $\rho_{\delta}$ is close to $\rho_{\rm ss}$.
@@
 $\mathcal{L}_{\phi}[\rho]=\gamma_{\phi}\sum_{i=1}^{N}\left(\sigma_{i}^{z}\rho\sigma_{i}^{z}-\rho\right),$ (29)
-which could be interpreted as a partial and local reset channel (only the $z$-axis is affected, and only local one-body jump operators are involoved).
+which could be interpreted as a partial and local reset channel (only the $z$-axis is affected, and only local one-body jump operators are involved).
 As shown in Fig.
 3, the relaxation is accelerated significantly.
-Figure 3: Dephasing noise induce acceleration of relaxation in a 5-site TFIM.
+Figure 3: Dephasing noise induces acceleration of relaxation in a 5-site TFIM.
 Parameters: (a) $J=1.0,\ g=2.0,\ \gamma=0.5,\ \beta=0.1J,\ t_{s}=0.2\tau_{2}$ (b) $J=1.0,\ g=2.0,\ \gamma=0.5,\ \beta=0.1J,\ t_{s}=0.8\tau_{2}$.
-A theoretical argument supporting the finding is provided in [59].